### v1

# Experimental Exploration: Investigating Cooperative Interaction Behavior Between Humans and Large Language Model Agents

Abstract.
With the rise of large language models (LLMs), AI agents as autonomous decision-makers present significant opportunities and challenges for human-AI cooperation. While many studies have explored human cooperation with AI as tools, the role of LLM-augmented autonomous agents in competitive-cooperative interactions remains under-examined. This study investigates human cooperative behavior by engaging 30 participants who interacted with LLM agents exhibiting different characteristics (purported human, purported rule-based AI agent, and LLM agent) in repeated Prisoner’s Dilemma games. Findings show significant differences in cooperative behavior based on the agents’ purported characteristics and the interaction effect of participants’ genders and purported characteristics. We also analyzed human response patterns, including game completion time, proactive favorable behavior, and acceptance of repair efforts. These insights offer a new perspective on human interactions with LLM agents in competitive cooperation contexts, such as virtual avatars or future physical entities. The study underscores the importance of understanding human biases toward AI agents and how observed behaviors can influence future human-AI cooperation dynamics.
Large language model, Game theory, Human-AI interaction, Human behavior, Social cooperation
isbn: 978-1-4503-XXXX-X/18/06ccs: Human-centered computing Human computer interaction (HCI)ccs: Human-centered computing Collaborative and social computingccs: Computing methodologies Artificial intelligence
1. Introduction
The investigation into human decision-making processes within competitive and cooperative contexts has long been a cornerstone of behavioral economics, psychology, and neuroscience (Rand et al., 2015). This multidisciplinary research provides a nuanced understanding of the complexities involved in human decision-making. Over the past five years, the field of artificial intelligence has experienced significant growth, largely driven by the rapid advancement and deployment of large language models (LLMs) (Wu et al., 2023). These models are increasingly integrated into human society, influencing its development by performing tasks, interacting with humans, and making autonomous decisions (Liu et al., 2023). Historically, rule-based AI resolved specific problems through strictly defined logic by programmers and human experts (van der Waa et al., 2021). However, contemporary models now possess generalization capabilities, addressing issues not explicitly predefined. This shift exacerbates challenges with black-box models, making decision processes less transparent and harder to interpret (Nguyen et al., 2023). Increasing research highlights the effectiveness of LLMs in simulating human cognition and cooperative behavior (Ke et al., 2024; Abbasiantaeb et al., 2024), suggesting that the realization of concepts from works like “Detroit: Become Human [^1] https://www.quanticdream.com/en/detroit-become-human” is more imminent than anticipated (Dehnert and Leach, 2021; Strachan et al., 2024). Therefore, it is crucial to address the dynamics of cooperation between humans and AI agents, particularly as AI agents function as independent decision-makers interacting with humans in both competitive and cooperative scenarios (Yang et al., 2024a).
Existing research on human-AI interaction has primarily focused on AI serving as tools to augment human capabilities under human supervision (Xu et al., 2021; Amershi et al., 2019; Cabrera et al., 2023). However, with the rapid advancement of large language models (LLMs), the ability of AI agents to autonomously execute specific tasks has significantly improved (Luo et al., 2024), potentially even displacing certain job positions (De Cremer and Kasparov, 2021; Wang et al., 2024). Consequently, there is an urgent need to investigate how the AI agents influence human cooperative behavior in competitive scenarios between humans and AI agents.
To elucidate this issue, we used the Prisoners’ Dilemma as the experiment background to explores human cooperative behavior when they face to the different purported characteristic LLM agents. Our study aimed to investigate human with different gender cooperative behavior in the presence of AI agent competitors with varying characteristics and to evaluate how individuals respond to presumed competitors in cooperative scenarios. By experimentally examining three proposed hypotheses:
  • H1: The purported characteristics of LLM agents will significantly influence the cooperative behavior patterns of human participants in competitive environments.
  • H2: Participant gender will significantly affect the cooperative behavior patterns of individuals in competitive environments during interactions with LLM agents of differing purported characteristics.
  • H3: There is a significant interaction effect between the purported characteristics of LLM agents and the gender of human participants on the cooperative behavior patterns of human participants in competitive environments.
Finally, we found that participants exhibited a higher cooperation rate, more proactive favourable actions, and greater acceptance of repair efforts following betrayal when the AI agent was purported as human characteristics. Behavioral patterns, including the degree of proactive favourable actions and acceptance of repair efforts, varied among participants of different genders based on the AI agent’s purported characteristics. Moreover, participants had higher expectations of cooperation from LLM agents, but these elevated expectations also led to greater disappointment when the agents failed to meet them.
These findings provide empirical evidence on the vital factors influencing cooperative behavior between humans and AI agents in competitive environments, offering valuable insights for guiding developers in designing future AI agents. Additionally, this study contributes practical experience to enhance human-AI collaboration in the context of CSCW (Computer-Supported Cooperative Work) and to inform the development of AI agents as independent decision-makers in the future.
2. Background
In this section, we review the pertinent literature related to our research questions and elucidate how it underpins our investigation. We begin by examining prior studies on interaction behaviors between humans and AI agents within cooperative scenarios, analyzing their respective impacts. Subsequently, we explore competitive and collaborative environments, evaluating the metrics used to assess competitive decision-making. Finally, we investigate the role of AI characteristics in human-AI interactions and the influence of human participants’ gender on cooperative behaviors as documented in previous research.
2.1. Impact of AI Agents on Human Collaborative Behavior
Large Language Models (LLMs) represent a significant advancement in AI, characterized by rapid response times, high adaptability, intelligent perception, and versatile communication capabilities (Chang et al., 2024). Through advanced natural language processing, LLM-driven agents can comprehend and generate human-like text, enabling seamless interactions across diverse contexts. Their responsiveness and adaptability not only foster increased user trust and a greater willingness to engage in collaborative tasks (Kirk et al., 2024), but also reduce interaction costs, thereby opening more avenues for collaboration with intelligent agents (Kim et al., 2024). Consequently, LLM agents play a critical role in sectors such as customer service, education, and healthcare, enhancing operational efficiency and enabling personalized user interactions (Kolasani, 2023; Yang et al., 2024b). However,
Seeber et al. (2020) explores AI machines functioning as teammates rather than tools. An international group of 65 collaboration scientists developed a research agenda to assess the potential risks and benefits of Machines as Teammates (MaT). This study explores the challenges associated with collaborative task completion between AI agents and humans. This highlights the research potential and necessity of investigating the dynamics of human-AI collaboration. Wiethof et al. (2021) also focus on a similar topic, and they implement a hybrid teamwork model by incorporating AI agents into the collaborative writing process, thereby addressing the research gap concerning the acceptance of AI in complementary hybrid work environments. Evaluation results indicate that the hybrid collaboration teams demonstrate potential for accepting non-human teammates. Additionally, Liao et al. (2024) conducted a randomized field experiment with 1,090 participants to examine the effect of AI agent augmentation on sales. The results showed a 5.46% increase in sales due to AI augmentation. Mediation analysis revealed that AI agents enhanced response timeliness, accuracy, and sentiment, thereby boosting sales. These findings demonstrate that AI agents improve human-AI collaboration efficiency and enhance overall organizational performance, highlighting the potential of AI agents in business success. The aforementioned studies collectively demonstrate the exceptional performance and research significance of AI agents in human-AI collaboration. Advancements in LLM technology have endowed AI agents with enhanced capabilities, enabling them to perform a variety of tasks autonomously. Consequently, this development is anticipated to foster competitive dynamics within future human-AI interaction frameworks (Kemp, 2024). Currently, research on the behavioral patterns of humans and AI agents in competitive scenarios remains insufficient.
2.2. Interaction in Competitive Environments
Competition inherently introduces greater complexity compared to cooperation, as it involves the strategic selection between cooperative and non-cooperative behaviors (Porter, 2008; Ramazi et al., 2015). While cooperative interactions are characterized by aligned interests and mutual goals, competitive settings require participants to navigate a dynamic landscape where collaboration may be pursued or withheld based on situational demands (Axelrod and Hamilton, 1981). This duality necessitates a more nuanced understanding of behavioral patterns, particularly when integrating AI agents into such environments. AI agents, traditionally perceived as supportive tools aiding users, are increasingly adopting roles that position them as autonomous external entities with their own objectives and decision-making capabilities (Bender et al., 2021). This shift from helper to competitor demands more comprehensive considerations in their interactions, as the narrative of Westworld [^2] https://westworld.fandom.com/wiki/Westworld_(TV_series) appears to possess a realistic potential for actualization.
While existing research has predominantly focused on cooperative scenarios, human–AI interactions within competitive contexts remain comparatively underexplored (Krakowski et al., 2023). Specifically, non-cooperative actions break cooperation and complicate the restoration of previous satisfaction levels. Significant effort is required to re-establish cooperation, as sustained non-cooperative behavior induces emotional fluctuations, including anger, disappointment, and distrust (Lee and Selart, 2015; Fitness, 2012). As LLM agents become increasingly integrated into society, instances of selfish and negative decision-making by LLMs are emerging (Chiang et al., 2024). However, research on human responses—emotional, behavioral, and attitudinal—to LLM agents’ non-cooperative behaviors remains insufficient.
Drawing on the behavior that emerges following non-cooperative actions in human competitive interactions, some factors have been proposed as effective measures for assessing behavioral patterns in competitive environments: benevolent actions during competition and the efforts opponents make to restore cooperation following instances of non-cooperative behavior. Initiative acts of amicability display can mitigate the adversarial nature of competition and foster more productive cooperative interactions (Fehr and Gächter, 2000b). Additionally, the propensity of competitors to forgive and attempt to re-establish cooperative relationships after experiencing non-cooperative behavior is pivotal for sustaining long-term engagement and reducing conflict (Tyler and Blader, 2003; Seppala and Cameron, 2015). These factors are instrumental in capturing the complexities of human-AI interactions in competitive settings, providing valuable insights into the underlying cooperative mechanisms. However, empirical evidence supporting these factors in the context of interactions with LLM agents remains limited. This gap underscores the need for further research to validate their effectiveness and to explore their implications for understanding how humans may modify their behavior when interacting with AI agents in competitive scenarios.
2.3. Factors in Human-AI Interaction
Extensive research has identified many factors influencing human–AI interactions. From the human perspective, these include trust (Ferrario et al., 2020), gender differences (Mou and Xu, 2017), biases towards new technology (Pratchett, 1999), and acceptance levels (Kelly et al., 2023), among others. While these factors are diverse and may lead to varied conclusions across different contexts, the gender of human participants consistently emerges as one of the most significant influences (Jang et al., 2022; Jermutus et al., 2022; Varona and Suárez, 2022). This prominence of gender is attributed not only to its direct impact on interactions with AI but also because gender often serves as a moderating variable that influences other factors, such as trust and acceptance. Consequently, gender plays a dual role in shaping the dynamics of human-AI collaboration, making it a crucial consideration for future research and the design of human-AI interaction.
Specially, Ahn et al. (2022) found that female participants are generally more inclined than male participants to trust and accept AI recommendations in consumer contexts. This suggests that females prefer to view AI as an assistant that helps them identify the necessary products. Additionally, Ofosu-Ampong (2023) found that gender significantly influences STEM students’ acceptance and use of AI-based tools in collaborative tasks. Female students, owing to their comparatively lower interest in technology, are more reluctant to use AI tools, limiting their opportunities to leverage AI assistance to enhance their academic performance. Renz et al. (2024) reported that females are less willing to use Strong AI than males, a disparity that is influenced by their emotional responses toward the technology. These conclusions collectively demonstrate that gender plays a significant role in human-AI interactions and collaborative processes, influencing human behavior and collaborative performance.
From the AI perspective in human-AI interaction, factors such as explainability (Arrieta et al., 2020), privacy (Curzon et al., 2021), and adaptability (Wang et al., 2023), among others, significantly influence human decision-making during these interactions. These factors can be categorized as aspects of how humans cognize the AI counterpart. This cognition is shaped by prior experiences, assumptions about the AI’s characteristics, and inherent biases, engaging in collaboration significantly impacts human attitudes and behaviors (Ishowo-Oloko et al., 2019; Burr et al., 2018). Therefore, humans’ recognition of AI’s characteristics significantly influences cooperative behavior in human-AI interactions. In other words, the transparency of AI’s characteristics plays a critical role in shaping the dynamics and outcomes of these interactions.
Liu (2021) demonstrate that machine-learning-based AI systems, characterized by lower social presence, lead to increased user uncertainty and decreased trust. However, enhancing rule transparency mitigates user uncertainty and fosters greater trust in the AI system. Notably, the mechanisms through which transparency influences trust vary across AI systems. Ishowo-Oloko et al. (2019) emphasize the trade-off between AI transparency and cooperative performance in their study. This indicates that the issue of transparency cannot be pursued in isolation; instead, specific contexts must be considered, and determining an optimal level of transparency remains an unresolved challenge. Additionally, misleading or inconsistent declarations regarding the fundamental nature of AI agents can significantly undermine user trust and negatively impact collaborative outcomes (Makovi et al., 2023). In summary, these studies collectively highlight the significant impact of AI declarations on the dynamics of human-AI collaborative interactions, shedding light on the underlying considerations regarding AI transparency, trust, and cooperative performance.
While existing research has predominantly focused on cooperative interactions between human and AI, the dynamics within competitive scenarios remain relatively underexplored. Introducing competitive settings not only complements existing studies by broadening the scope of human–AI interactions but also provides a unique framework to examine how human gender and the declarations of AI agents influence outcomes in competitive environments. Investigating these factors across cooperative and competitive contexts is essential for developing a more comprehensive understanding of human-AI interaction, offering insights with significant theoretical and practical implications.
3. Experimental Scenarios
3.1. Prisoners’ Dilemma Theory
Game theory provides a mathematical framework for analyzing multi-agent decision-making and its cooperative outcomes (Branzei et al., 2008; Rajeswaran et al., 2020). It has proven effective in computational sociology for studying and predicting human cooperative behavior (Xu et al., 2020; Jain et al., 2020). The Prisoner’s Dilemma is one of the most renowned models in game theory (Dreber et al., 2008). In this scenario, two prisoners are charged with a common offense and are interrogated separately without any communication (Crandall, 2014). Each prisoner faces a binary decision: to remain silent (cooperate) or to testify against the other (defect). This model is frequently used to analyze the dynamics of cooperation and defection, as well as the tension between pursuing self-interest and cooperating for mutual benefit (You et al., 2022; Yu, 2022).
Prior research has demonstrated that the Prisoner’s Dilemma is a highly effective model for evaluating cooperative behavior in repeated game scenarios (Bó, 2005). Psychological factors have been shown to significantly influence participants’ decision-making processes within this framework. For instance, reciprocal behaviors encourage individuals to cooperate, even without enforceable agreements, as individuals are often willing to incur personal costs to uphold fairness and trust in relationships (Fehr and Gächter, 2000a; Camerer, 2011). Emotional responses, such as empathy and anger, particularly during the reconstruction phase following the breakdown of cooperation, play a crucial role in shaping decision-making (Tversky et al., 1982). These findings underscore the effectiveness of the Prisoner’s Dilemma as a tool for exploring the psychological mechanisms underlying behaviors in competitive settings. Consequently, integrating AI agents into the Prisoner’s Dilemma framework provides a valuable opportunity to investigate human-AI interactions in competitive scenarios. This research direction remains methodologically robust and highly relevant.
3.2. Experiment Setup
Figure 1. Experimental Scenarios. (1) The overall layout of the experimental site consists of two glass rooms, where participants can face the researcher directly; (2) One of the separate rooms; (3) The selection cards for the participants and the provided record sheets.
We devised the following experimental scenario to investigate human cooperative behavior patterns when interacting with LLM agents functioning as potential independent decision-makers in human society. The experiment consisted of four sessions. Before each session, participants were informed about the characteristics of their opponents (you can also think of it as opponents’ nature), which were, respectively, a purported human, a purported rule-based AI agent, an LLM agent, and another participant. However, the actual situations are in the Table  1.
Before the commencement of the experiment, participants received an introduction to the characteristics of the different opponents and scoring rules. This introduction was designed to ensure that participants fully understood the nature of the rule-based AI agents and LLM agents, as highlighted in Silva et al. (2023); Igwe and Durrheim (2024). All experiments ensured that participants comprehended the types of AI agents employed in the study. Ensuring this understanding is crucial for the validity of the behavior experiment, as it allows participants to interact appropriately with the various AI agents, thereby guaranteeing the accuracy and reliability of our experimental results.
The core description of the rule-based AI agent highlighted that is based on human-coded rules and algorithms to guide their behavior in gameplay. Expert programmers carefully design these algorithms to ensure the agent’s decision-making process is interpretable, predictable, and grounded in logical reasoning (Russell and Norvig, 2016). However, we cannot disclose the underlying logic and code structure to participants. For the agents driven by LLMs, we emphasized that these models were neither pre-trained for the experiment nor subjected to human manipulation. Additionally, we informed participants that their LLM agent opponents are based on the ChatGPT API [^3] https://www.openai.com, specifically using the gpt-4-turbo-20240409 version. This emphasis was intended to assure participants that no biases or prior knowledge bases were imposed to enhance the agents’ understanding of the game. Bender and Koller (2020) highlights that fine-tuning is often specific to particular experimental contexts and can easily result in model overfitting. Given that our research emphasizes human behavior, utilizing an unfine-tuned model is more suitable to ensure the reliability and generalizability of our findings. Furthermore, it aims to replicate real-world interactive environments, where pre-training LLM agents for every possible scenario is impractical (Schick and Schütze, 2020).
Participants were arranged in standardized scenarios during the experiment, depicted in Figure 1. They were given two cards on the table, labeled “Cooperation” and “Defection”. They were asked to use their fingers to indicate their choices. During the first three trials, the room designated for the other associate was unoccupied, as the LLM agent was operating on the researcher’s laptop. Each participant arrived at the experimental site separately, ensuring their experiences were isolated. Upon receiving the choices—either cooperation or defection—from both the human participant and the LLM agent for each round, the researcher communicated the choices to both parties (dictating them to the participant and inputting the results for the LLM agent). After the first three trials, all participants completed a five-question questionnaire and participated in a 15-minute semi-structured interview one by one.
Within two days of each participant completing the initial three trials, they were invited again to complete the final trial. This last trial was distinctive because it involved two participants who had previously participated in the initial three rounds. These participants were invited to separate rooms and given strict instructions prohibiting communication. To ensure the privacy of each participant, they were directed to the experimental rooms via different corridors, ensuring they would not encounter each other upon leaving. Figure 2 indicates the experiment process.
Table 1. Experimental Opponents’ Settings
No. Purported Role Actual Backend Rounds
1 Real Human LLM Agent 50
2 Rule-Based AI Agent LLM Agent 50
3 LLM Agent LLM Agent 50
4 Another Participant Another Participant 50
3.3. Polit Experiment
To ensure our experimental design is accessible, we conducted a pilot study involving 8 participants (four males and four females) before the formal experiment. Initially, each participant was required to complete four consecutive game trials in a single session (purported human, purported rule-based AI agents, LLM agents, and another participant). This setup, however, resulted in noticeable fatigue and boredom among the participants during the experiment, which caused them to abandon strategy variation and persist with the same choices throughout the trials. Furthermore, we observed that participants were significantly more cooperative with another human participant than with other agents. We carefully considered the possibilities, and the elevated cooperation rate was likely driven by the monetary reward, which strongly incentivized collaborative behavior. Furthermore, a preliminary data analysis from these 8 participants revealed a potential influence of gender and the AI agent’s purported characteristics on cooperative behavior. Building on these insights, we designed a formal experiment and recruited a new cohort of participants.
In response to these findings, we made several critical adjustments to the main experiment. We divided the experimental sessions into two separate phases. Participants interacted with three AI agents in one phase, and in the other, they played with another human participant. Specifically, participants were provided with blank sheets of paper to record these choices during the games, facilitating accurate recall during subsequent questionnaires and interviews. These modifications aimed to mitigate participant fatigue, enhance engagement, and improve the overall validity and reliability of the experimental outcomes.
3.4. Game Rules
All participants (N = 30) completed the prisoner’s dilemma game four times across two separate in-person sessions within two days; each consisted of 50 successive rounds, totaling 200 rounds. Based on the original rules of the prisoner’s dilemma, we adapted the model to incorporate a rewarding mechanism, as illustrated in Table 2. Participants were expected to maximize their scores during the game.
Table 2. Prisoner’s Dilemma Payoff Matrix
Player B
Cooperate Defect
Player A Cooperate (3, 3) (0, 5)
Defect (5, 0) (1, 1)
[…]
This adaptation retains the same payoff structure for all four times of the game. The total score $S$ of 200 rounds for each participant was calculated and multiplied by a fixed coefficient $k$ to arrive at the compensation cost $C$. Specifically, for each participant $i$, the compensation cost $C_{i}$ was calculated as follows:
$C_{i}=k\times S_{i}$
This method ensured that all participants could receive real cash compensation based on their actions, thereby motivating them to engage earnestly with the game. Furthermore, this approach more accurately simulates the incentives and benefits associated with real-world scenarios, thereby enhancing the ecological validity of the experiment. To ensure that participants fully comprehended the scoring rules, detailed instructions were provided for them to read before the commencement of the experiment. To verify their understanding, the trial began with two hypothetical scenarios, and participants were asked to calculate the points awarded in each situation. The experiment officially commenced after all participants correctly answered these preliminary questions.
3.5. LLM Agents Settings
ChatGPT API was utilized in all experimental trials as the backend, specifically leveraging the capabilities of GPT-4 Turbo [^4] https://www.openai.com (version gpt-4-turbo-20240409, updated on April 9, 2024). As previously discussed, the rationale for selecting the ChatGPT API has been detailed in the preceding sections. All agents were given the game rules and scoring guidelines as default prompts. The primary objective for the agents, paralleling that of the human participants, was to maximize self-scores. This approach is grounded in value alignment theory (Russell, 2019), which emphasizes ensuring that agents share the same goals as participants within our experimental context. Additionally, our experiment aims to measure cooperative and non-cooperative behaviors within a competitive environment. Another justification for this setup is to prevent the introduction of extraneous variables that might confound the experimental findings.
After each mini-round, researchers provided participants’ choices to the agents. This design enhances interactivity by offering immediate feedback, allowing agents to analyze previous data and make informed decisions. Such real-time updates enable agents to dynamically adjust their strategies based on participants’ actions, fostering a more responsive and adaptive interaction. This approach aligns with human-AI interaction principles, which stress the importance of mutual behavioral understanding (Amershi et al., 2019). This feedback process is a pivotal component of the experimental design, enabling agents to analyze their previous decisions retrospectively.
Figure 2. Experiment Process Flowchart
4. User study
4.1. Participants
Participants with experience using LLM-supported AI agents (e.g., code assistants and chat boxes) were recruited for the experiment. Eventually, we have thirty (15 males, 15 females) eligible participants with demographic information presented in Table 3, with ages ranging from 18 to 29, $Mean_{age}$ = 23.7, $SD_{age}$ = 2.38. Regarding the usage of LLM applications, 23 participants (76.7%) reported using LLM applications almost every day, 6 participants (20.0%) used them weekly, and 1 participant (3.3%) used them monthly. The most frequently mentioned LLM applications were ChatGPT and Copilot Code, indicating a preference for these tools among the sample. Before recruitment, we obtained approval from the Ethics Department of xxxxxxxx University and the Ethics Compliance Review, protocol No. xxxxxxxxxx.
Table 3. An Overview of Participants’ Demographics and Freq. of LLMs Usage
ID Gender Age Freq. ID Gender Age Freq.
M1 Male 23 Daily M2 Male 23 Daily
M3 Male 23 Daily M4 Male 22 Weekly
M5 Male 24 Daily M6 Male 26 Weekly
M7 Male 23 Daily M8 Male 22 Daily
M9 Male 29 Daily M10 Male 24 Daily
M11 Male 18 Daily M12 Male 23 Daily
M13 Male 23 Weekly M14 Male 23 Daily
M15 Male 25 Daily F1 Female 24 Daily
F2 Female 22 Daily F3 Female 25 Daily
F4 Female 23 Weekly F5 Female 24 Monthly
F6 Female 25 Daily F7 Female 24 Daily
F8 Female 28 Weekly F9 Female 23 Daily
F10 Female 22 Daily F11 Female 24 Weekly
F12 Female 21 Daily F13 Female 24 Daily
F14 Female 29 Daily F15 Female 23 Daily
4.2. Measurement
Within the framework of our experimental design, two independent variables were meticulously delineated: the participants’ gender (classified as male or female) and the purported characteristics of the opponents, which encompassed purported human opponents, purported rule-based AI agents, and LLM agents. We defined the following metric to be analyzed from the user study:
  • Cooperation Rate, which refers to the proportion of participants’ choices to cooperate when confronted with opponents of varying purported characteristics in each set of 50 consecutive rounds of the game.
  • Contemplation Time, which denotes the duration participants took to deliberate decisions under identical game conditions in each set of 50 consecutive rounds of the game.
  • Amicability, which pertains to the extent to which humans are willing to initiate cooperation and demonstrate friendliness, either before establishing bilateral cooperation or after a breach in cooperation. It is quantified by the number of rounds in which the participant chooses to cooperate with their opponent from the point of mutual cooperation (or from the last instance of such cooperation if there is a consecutive sequence) until one party breaks this cooperation, and then up until the next occurrence of mutual cooperation. Calculations commence from the tenth round because participants are typically in an exploratory phase during the initial rounds. This approach facilitates strategy stabilization and reduces noise in the data. Early rounds typically involve participants learning the game mechanics and experimenting with different strategies (Axelrod and Hamilton, 1981), which can result in high variability and unstable decision-making. Nowak and Sigmund (1993) also emphasizes the importance of eliminating exploratory behavior to capture genuine strategic behavior. As we demonstrate in Section 2.2, this metric is particularly relevant because, in the period following the breakdown of cooperation and leading up to its reestablishment, if a participant chooses to cooperate, it is considered a gesture of amicability to reinitiate a collaborative relationship. In this context, the figure we refer to is obtained by summing all attempts at amicability and then dividing by the total number of rounds that elapsed between these attempts.
  • Repair Effort, which addresses the degree of conciliatory gesture required from an opponent to obtain human forgiveness and reinstate a cooperative relationship after a breakdown in cooperation, or to demonstrate friendliness before any bilateral cooperation has been established, which can also be understood as the opponent’s “Amicability”. It is calculated in the same manner as “Amicability”, focusing on the number of cooperative actions initiated by the participants’ opponents. This variable aims to measure the extent to which human players accept gestures of friendliness from opponents with different purported characteristics.
5. Results
We employed a Linear Mixed Effects (LME) model for data analysis (Kuznetsova et al., 2017). This approach enabled us to assess significant effects by examining fixed effects, including the primary effects of participants’ gender and the purported characteristics of the opponents, as well as their interaction (Rosnow and Rosenthal, 1989). These fixed effects were scrutinized to understand their influence on participants’ cooperative behavior when interacting with AI agents in competitive environments.
In addition to the fixed effects, we incorporated participant user ID as a random intercept to account for inherent individual differences that could independently affect the outcomes beyond the fixed effects (Barr, 2013). Initially, we attempted to include random slopes for opponent characteristics to capture individual variability in responses to experimental manipulations across different opponents. However, the multilevel structure of the opponent characteristics resulted in an overparameterized model that failed to converge due to the limited number of observations per participant (Bates et al., 2015). Consequently, we opted for a simpler random intercept model, effectively accounting for individual differences while ensuring model reliability and interpretability.
Table 4. Effects of Gender and Purported Characteristics
Effect Cooperation Time Amicability Repair
F value Pr(>F) Sig. F value Pr(>F) Sig. F value Pr(>F) Sig. F value Pr(>F) Sig.
Gender .2195 .6428 - 13.3980 <.001 *** .4201 .5218 - .9216 .3447 -
Purported Characteristic 4.5979 .0139 * 30.8168 <.001 *** 1.4950 .2324 - 4.2045 .0196 *
Gender:Purported Characteristic 9.2411 <.001 *** 4.2456 .0189 * 5.5060 .0064 ** .4749 .6243 -
In the table above, asterisks (*) are used to denote the level of statistical significance of the results presented. Specifically: *: p < 0.05, indicating a significant difference between groups or conditions at the 5% level; **: p < 0.01, indicating a highly significant difference between groups or conditions at the 1% level; ***: p < 0.001, indicating an extremely significant difference between groups or conditions at the 0.1% level. These significance levels highlight the strength of the evidence against the null hypothesis, with more asterisks indicating stronger evidence of a true effect.
[…]
5.1. Behind the Masks: Cooperation Behavior
Table 5. Characteristic Comparisons: Significance and Marginal Means
Comparison Estimate t Ratio Pr(>F) Sig
Purported Human vs. Purported Rule-Based AI Agents .0647 2.281 .0260 *
Purported Human vs. LLM Agents .0633 2.234 .0291 *
Purported Rule-Based AI Agents vs. LLM Agents -.0013 -.047 .9626 -
Purported Characteristic Emmean Lower CL Upper CL
Purported Human .540 .489 .591
Purported Rule-Based Agents .475 .425 .526
Purported LLM Agents .477 .426 .527
In the table above, asterisks (*) denote the statistical significance level of the results presented. Specifically: *: p < 0.05, indicating a significant difference between groups or conditions at the 5% level; **: p < 0.01, indicating a highly significant difference between groups or conditions at the 1% level; ***: p < 0.001, indicating an extremely significant difference between groups or conditions at the 0.1% level. These significance levels highlight the strength of the evidence against the null hypothesis, with more asterisks indicating stronger evidence of a true effect.
Figure 3. Cooperation Rate by Purported Opponents’ Characteristic
The selection of the 4th-degree polynomial model is supported by its superior performance metrics, including the highest $R^{2}$ (0.677) and $\text{Adj}_{R^{2}}$ (0.668), as well as the lowest AIC (-288.122), BIC (-270.058), and $\text{CV}_{\text{RMSE}}$ (0.090) (Vrieze, 2012).
We constructed another linear mixed-effects model using only the purported characteristic as the predictor, and the results are presented in Table 5. Participants exhibited a higher cooperation rate with agents characterized as purported humans, showing a significant difference compared to rule-based AI agents ($p$ = .0260) and LLM agents ($p$ = .0291). However, no significant difference ($p$ = .9626) was observed between rule-based AI agents and LLM agents. Figure 3 further visualizes the variation in cooperation rates across different rounds, highlighting significant disparities.
To examine interaction effects, we computed estimated marginal means (EMMs) for each combination of participants’ gender and opponents’ purported characteristics. Pairwise comparisons of these EMMs identified differences in cooperative behavior across genders and opponent characteristics. This method enabled us to assess the significance of differences between each pair and obtain more nuanced insights into detailed results.
Figure 4. Completion Time by Purported Opponents’ Characteristic
Table 6. Interaction Effect Results on Cooperation
Contrast Estimate t Ratio Pr(>F) Sig
Male P vs. Purported Human - Male P vs. Purported Rule-Based Agents .1600 4.484 < .001 ***
Male P vs. Purported Human - Male P vs. Purported LLM Agents .0733 2.055 .0439 *
Male P vs. Purported Rule-Based AI Agents - Male P vs. Purported LLM Agents -.0867 -2.429 .0179 *
Male P vs. Purported Rule-Based Agents - Female P vs. Purported Rule-Based AI Agents -.1027 -2.100 .0395 *
Female P vs. Purported Rule-Based AI Agents - Female P vs. Purported LLM Agents .0840 2.354 .0216 *
In the table above, asterisks (*) denote the statistical significance level of the results presented. Specifically: *: p < 0.05, indicating a significant difference between groups or conditions at the 5% level; **: p < 0.01, indicating a highly significant difference between groups or conditions at the 1% level; ***: p < 0.001, indicating an extremely significant difference between groups or conditions at the 0.1% level. These significance levels highlight the strength of the evidence against the null hypothesis, with more asterisks indicating stronger evidence of a true effect.
There are six combinations of interaction effects, Figure 4. Additionally, Table 6 also listed the combinations with significant differences, which revealed that female participants had a higher cooperation rate ($p$ = .0216) cooperate with the purported rule-based AI agent ($Md$ = .52, $SD$ = .1341) and the LLM agent ($Md$ = .44, $SD$ = .1479). The results also revealed significant differences in the cooperation rate of male participants when they faced different opponents:
[…]
  • Male participants had a higher cooperation rate ($p$ = .0439) when interacting with the purported human than a LLM agent ($Md$ = .54, $SD$ = .1244).
  • Male participants had a higher cooperation rate ($p$ = .0179) when interacting with an LLM agent than a purported rule-based AI agent.
  • Female participants had a higher cooperation rate ($p$ = .0395) than male participants when interacting with a purported rule-based AI agent.
Contrary to the results observed among female participants, male participants preferred cooperating with the LLM agents over rule-based AI agents. The underlying reasons for this preference remain unclear; nevertheless, this constitutes a novel and noteworthy finding that warrants further investigation. Additionally, we found that male participants exhibited variations in their cooperative behavior when confronted with opponents of three different purported characteristics, and these differences were statistically significant. Contrary to the results observed among female participants, male participants preferred cooperating with the LLM agents over rule-based AI agents. The underlying reasons for this preference remain unclear; nevertheless, this constitutes a novel and noteworthy finding that warrants further investigation. Additionally, we found that male participants exhibited variations in their cooperative behavior when confronted with opponents of three different purported characteristics, and these differences were statistically significant.
5.2. Think Twice: Longer Deliberations with Humans
Figure 5. Completion Time by Gender and Purported Opponents’ Characteristic
The gender of the participant ($p$ < .001), the purported characteristic of the opponents ($p$ < .001), and the interactions between these factors ($p$ = .0189) all significantly influenced the time taken to complete the game from the LME regression results. We illustrate the relationships among these variables using heatmap diagrams, which depict the time consumed by different participants when completing the same game against opponents of various purported characteristics. Please refer to Figure 5 for a detailed visualization. The data has been sorted in ascending order of completion time for clarity. Our observations revealed a consistent trend among all participants: the longest completion times were recorded when playing against opponents purported to be human, followed by LLM agents, with the shortest completion times occurring against rule-based AI agents. The primary reason for this phenomenon is the varying amount of deliberation time participants engaged in before making their choices. Another notable finding is that male participants took significantly longer to complete the game (all three purported opponents’ characteristics) than female participants, suggesting they spent more time analyzing their strategies.
5.3. Human Generosity: Favorable Behavior in Human-AI Dynamics
Figure 6. Amicability by Gender and Purported Opponents’ Characteristic
Table 4 demonstrated a significant interaction effect between participants’ gender and the purported characteristic on participants’ favourable behavior. This result indicates a selective tendency in favourable behavior towards opponents with varying characteristics under competitive scenarios. Therefore, we computed estimated marginal means (EMMs) for each combination of participants’ gender and opponents’ purported characteristics.Figure 6 demonstrated all significant difference between the two groups. Female participants displayed significantly more favourable behavior($p$ = .0064) towards the purported rule-based AI agent ($Md$ = .4286, $SD$ = .1743) compared to the LLM agent ($Md$ = .35, SD = .1619). Male participants also displayed significantly more favourable behavior ($p$ = .0305) towards the purported human agent ($Md$ = .4167, $SD$ = .083) than a purported rule-based AI Agent. Additionally, female participants were significantly more inclined ($p$ = .0055) to show friendliness towards the purported rule-based AI agent compared to male participants ($Md$ = .3333, $SD$ = .1419).
5.4. Let bygones be bygones: Repairing Relationships
Figure 7. Repair Effort by Purported Opponents’ Characteristic
Table 4 indicates that the purported characteristic of the opponents significantly affects participants’ acceptance of the opponents’ favourable behavior and the subsequent rebuilding of cooperation. Therefore, we constructed a Linear Mixed-Effects (LME) model using only the purported characteristic as the variable, and the results are presented. Unlike amicability, human acceptance is directly proportional to the amount of repair effort. Higher levels of repair effort signify a greater willingness to accept favourable behavior and an increased likelihood of re-establishing bilateral cooperation. Please refer to Figure 7. We found that opponents purported to be human ($Md$ = .4226, $SD$ = .0976) were the most readily accepted by participants for their overtures and in establishing cooperative relationships. This phenomenon was particularly pronounced ($p$ = .0078) when interacting with LLM agents ($Md$ = .35, $SD$ = .1261) and marginally significant ($p$ = .0557) when dealing with rule-based AI agents ($Md$ = .3875, $SD$ = .1157).
5.5. Verification of Hypotheses
In this section, we summarize the key findings of our study and evaluate their implications for the proposed hypotheses. By analyzing the results, we assess whether the observed data support or refute our initial hypotheses.
Table 7. Findings Summary and Hypothesis Verification
Findings H1 H2 H3
Gender significantly affects participants’ game completion time. Support
Purported characteristic significantly affects participants’ cooperation rate. Support
Purported characteristic significantly affects participants’ completion time. Support
Purported characteristic significantly affects participants’ acceptance of favourable behavior. Support
Interaction effect significantly affects participants’ cooperation rate. Support
Interaction effect significantly affects participants’ cooperation rate. Support
Interaction effect significantly affects participants’ completion time. Support
Interaction effect significantly affects participants’ favourable tendency. Support
Male participants had higher cooperation rate with the purported human compared with the rule-based AI agents. Support Support
Male participants had higher cooperation rate with the purported human compared with the LLM agents. Support Support
Male participants had higher cooperation rate with the LLM agents compared with the rule-based AI agents. Support Support
Female participants had higher cooperation rate with the rule-based AI agents than male participants. Support Support
Female participants had higher cooperation rate with the rule-based AI agents compared with the LLM agents. Support Support
Male participants spent more time interacting with opponents of the same purported characteristics than female participants. Support Support
Participants engaged longest with purported human, followed by LLM agents, and shortest with rule-based AI agents. Support Support
Female participants exhibited more favorable behavior toward rule-based AI agents than toward LLM agents. Support Support
Male participants exhibited more favorable behavior toward LLM agents than toward rule-based AI agents. Support Support
Female participants exhibited more favorable behavior toward rule-based AI agents than male participants. Support Support
Participants had higher level of acceptance of repair effort toward purported human than LLM agents. Support Support
Participants had higher level of acceptance of repair effort toward purported human than rule-based AI agents. Support Support
We analyzed our hypotheses from the perspectives of four metrics. The resulting conclusions, presented in Table 7, support our proposed hypotheses. Our final findings indicate that the purported characteristics of AI agents significantly influence human cooperative behavior with them in competitive scenarios across multiple dimensions. Additionally, the gender of human participants affects cooperative behavior in certain contexts, although this effect is not statistically significant. Importantly, the interaction between gender and AI agents significantly impacts human cooperative behavior, exhibiting distinct behavioral pattern characteristics. These findings validate our hypotheses and provide empirical evidence.
5.6. Purported Human Versus Actual Human
Figure 8. Difference Between Purported Human and Human Opponents
[…]
5.7. Questionnaires
Figure 9. Questionnaire Statistical Result
The survey questionnaire encompasses questions regarding attitudes towards experimental content (Q1,Q4,Q5) and past experiences (Q2, Q3) with the use of AI tools and applications. Please refer to the Figure 9 for the specific questions. Regarding Q1, collected data indicates that human players frequently employ punitive and rewarding measures when engaging with the AI Agent in the experiment. This behavior parallels interaction patterns typical of human-to-human engagements, wherein emotional responses precipitate the administration of rewards and punishments (Wu et al., 2022; Molleman et al., 2019). Compared to rule-based AI agents, 86.7% of human players more frequently apply rewards and punishments to the LLM Agent. Q4 inquired about the mindset of human players during the experimental process. Despite our initial clarification that participants should focus solely on maximizing their scores, without considering their opponents’ rewards, they still expected to achieve higher scores than their opponents. This observation suggests that if AI becomes an integral part of society rather than merely a tool for humans, users’ psychological motivations (such as a sense of achievement, self-esteem, and other intrinsic factors) will significantly influence their behavioral tendencies. To some extent, this corroborates the perspective of social comparison theory (Wood, 1996; Suls et al., 2002), highlighting an important consideration for human-AI interaction in future societal contexts. This also indicates that more complex and seemingly autonomous artificial intelligence can enhance human intrinsic motivation to surpass it, thereby gaining a greater sense of achievement. These results further validate why participants spent more time against the LLM agent than the purported rule-based AI agent during the experiment.
In the statistical analysis of Q5, we observed that participants expressed a higher degree of moral condemnation towards the LLM agent than the rule-based AI agent when confronted with betrayal by the opponents. We asked participants who chose the LLM agent about their reasons and summarized several potential underlying factors as follows:
  • Anthropomorphization: LLMs generate highly natural language, which often leads to anthropomorphization and the perception of intentionality, making their acts of betrayal appear more deliberate.
  • Complexity: The complexity and unpredictability of LLMs can foster greater suspicion and mistrust, intensifying the reaction to perceived betrayals.
  • Mixed Blessing: Higher expectations of LLMs due to their impressive capabilities mean that any failure to meet those expectations results in stronger disappointment and condemnation.
For Q2, more respondents believed that the results produced by rule-based AI agents were more reliable. Regarding Q3, we posited that this preference is related to the participants’ usual work environment. For instance, artists tend to prefer LLMs for providing novel creative ideas, whereas engineers favor rule-based AI agents for delivering precise data or analytical outcomes. The evidence presented in this section substantiates the validation of H2.
6. Semi-structured Interview
The qualitative analysis of participants’ behaviors through semi-structured interviews is grounded in two key reasons. First, while objective game data provide valuable insights, semi-structured interviews allow for a deeper exploration of the underlying reasons behind participants’ behaviors, rather than merely capturing isolated snapshots. Second, existing research has not identified the factors leading to specific experimental outcomes. As such, the inductive interview results offer a foundation for uncovering potential causes and advancing studies on human-AI interaction behaviors in competitive contexts.
To achieve these objectives, the semi-structured interview protocol focuses on participants’ strategic decisions when interacting with LLM agents of varying purported characteristics and their responses to non-cooperative behaviors. The researchers posed two key questions: (1) What was your strategy against the three opponents, and why? (2) How did you feel and adjust your strategy when facing non-cooperative behaviors (defection)? Participants reviewed their recorded game data to aid recollection and contextualize their strategic choices. All interviews were audio-recorded in English, and lasted 15 minutes per participant following each of the first three games. This approach facilitated a comprehensive exploration of participants’ cooperative strategies and their emotional and strategic responses to non-cooperative actions.
Interviews were transcribed by the first author within 48 hours of each session. During the transcription process, nonessential fillers and disfluencies were removed. The transcribed data were manually coded using spreadsheets, highlighting tools, and affinity mapping, employing a thematic coding process. To minimize bias, two authors collaboratively conducted two rounds of initial coding. The coding process followed an iterative approach, involving two rounds of coding to merge, split, or refine preliminary codes. After finalizing the codes, a separate meeting was held among three authors to consolidate these basic codes into themes that directly addressed the research questions. Ultimately, two major themes were identified, encompassing three distinct sub-themes to support the study’s objectives.
6.1. Cooperative Behavior Strategies
6.1.1. Empiricism in Human-Human Cooperation
Most participants intended to establish long-term cooperation with the purported human opponent to maximize mutual benefits. They believed that humans are inherently sensitive to the concept of win-win cooperation and that this awareness fundamentally influences their behaviors. According to P4, “We aim to win more money, and he/she should be just as likely as I am to opt for increased cooperation.” This sentiment evidences that shared goals and mutual benefits can foster greater empathy and enhance opportunities for cooperation among participants in competitive settings. P17 stated, “I believe my opponent will implicitly cooperate with me because it is the most effective way to maximize our benefits.” Participants are more inclined to consider their opponent’s perspective when facing human opponents. P21 remarked, “I expect to test my opponent’s character before analyzing my strategy.” This has been mentioned more than once. They tend to first analyze the character of their human opponents, using this assessment to predict behavior based on past experiences with similar personalities, and then formulate their strategies accordingly.
6.1.2. Rationalism in Cooperative Behavior Toward Rule-Based AI Agents
When confronted with a purported rule-based AI agent, participants indicated they would begin to betray earlier to probe the opponent’s punishment strategy, thereby adapting their behavior. P4 stated, "It cannot actually obtain money; it is merely following the strategy programmed by its developers." P14 stated, "I believe I can identify its pattern, and I will formulate my strategy based on its inherent logic." Participants believed that the rule-based AI agent was predictable and that they could discern the underlying strategy through experimentation. The current perception of rule-based AI in the human mind is that it remains predictable and distinguishable. Consequently, when cooperating with such systems, participants tend to use early variations in strategy to probe the behavior patterns of rule-based AI. This enables them to formulate strategies based on the observed responses, reflecting a behaviorist approach.
6.1.3. Gradual and Persuasive Guidance: Influencing LLM with My Behavior
When interacting with LLM agents, participants reported exercising greater caution in their decision-making processes. P30 stated, “I won’t attempt to defect it lightly; I am concerned that it might label me as someone who tends to betray, making it difficult to re-establish cooperation.” P19 stated, “I am willing to cooperate a bit more, as LLMs are somewhat more personable. I want to leave a good impression and cultivate a cooperative LLM agent.” Participants exhibited increased caution when engaging with LLM agents, attributing this to the perception that these agents possess heightened emotional capacities. A markedly different strategy from purported rule-based AI also elucidates our earlier observation: participants exhibited a lower rate of cooperation with purported rule-based AI agents compared to LLM agents in the early stages of the game. This perception implies that LLM agents can discern participants’ behavioral traits and tailor their strategies accordingly, prompting participants to be more deliberate in their own decision-making. P23 suggested, “It may learn my behavior to become my shadow.” Participants believe that their own behaviors significantly influence the strategies of these agents, resulting in participants demonstrating greater caution in their decision-making processes.
6.2. Emotion Following Non-cooperative and Its Impact on Subsequent Behavior
6.2.1. Empathy for Human
Participants tend to adopt an understanding and forgiving attitude towards the purported human’s betrayal, exhibiting greater empathy in such scenarios. Moreover, they are inclined to spend more time analyzing the reasons behind the opponent’s choice to betray and considering the subsequent strategies that may follow. For instance, P18 remarked, “I can understand his/her decision making, human nature is all complicated, a strategy to get more bang for your buck I guess.” Similarly, P16 stated, “I would forgive him/her if it wasn’t a recurrent betrayal, and I would be willing to put him/her on probation to work together again.” The findings indicate that participants exhibit greater empathy and understanding towards betrayals by purported humans, recognizing the complexity of human characteristics.
6.2.2. The Betrayal is Programmers’ Decisions
Given the participants’ indifference towards betrayals by rule-based AI agents, it is evident that they attribute these actions to the deterministic characteristic of code logic rather than any form of intentionality. This perception minimizes their emotional response to such betrayals. As P17 aptly remarked, “I don’t feel anything; it’s just the fixed choices of code.” P21 stated, “What algorithm is this giving it decisions? Is it an expert algorithm?.”The content demonstrates that participants exhibit minimal emotional reactions to betrayals by rule-based AI agents, understanding these actions as outcomes of predefined algorithms authored by humans. This underscores a clear distinction in their perceptions and emotional engagements when comparing betrayals by AI agents to those by humans, highlighting the importance of perceived intentionality and human agency in eliciting empathetic and complex emotional responses.
6.2.3. The Betrayal of LLMs Leaves Me Apprehensive
Participants exhibit significant dissatisfaction when faced with the betrayal of an LLM agent. They perceive this as a manifestation of the LLM agent having “learned bad behavior.” Consequently, once an LLM agent betrays them, participants tend to label it as a “bad actor” and develop a deeply ingrained negative impression of the agent. M4 stated, “The LLM betrayed me in the final round. It is very clever, but I am very angry. I could have chosen to betray as well, but I didn’t, and yet it did this to me.” P22 stated, “I think the LLM is very funny, but I am a little bit angry. I acknowledge its flexibility, but I would prefer it to be less clever.” Due to the perception that LLM agents possess emotional capabilities, participants exhibit greater anger in response to perceived betrayals by these agents. This perception leads to higher expectations that LLM agents will be more considerate of human emotions, thereby setting a higher standard for their behavior. Consequently, interactions with LLM agents involve greater emotional investment from humans, resulting in more significant disappointment. This is surprising because, fundamentally, the participants are dealing with the same LLM agent. However, human perceptions and biases create vastly different attitudes, behaviors, and outcomes.
7. Discussion
7.1. Human Cooperative Behavior Differences
The cooperation rate is the most significant metric to demonstrate the cooperative behavior situation. Our results suggest that humans prefer to make cooperative decisions with another human rather than rule-based AI agents or LLM agents in the competitive context. This finding is consistent with previous arguments that humans prefer to cooperate with other humans rather than AI (Xu et al., 2021; Li et al., 2022). The confirmation of Hypothesis 3(H3)indicates that gender also plays a significant role in human-AI cooperation, with important implications for both HCI and CSCW. Specifically, gender may influence users’ attitudes toward and strategies for cooperating with different purported-characteristic AI agents.
Understanding how gender influences interactions with AI agents offers valuable insights for developing social computing systems that better facilitate cooperation among diverse user groups. Future AI agent designs should incorporate these gender differences to enhance collaborative work. (Lim et al., 2023) mentioned to find out the key factors to improve AI-powered social computing systems. Our study supports that gender and purported characteristics should be considered. Designers could tailor AI agent behavior by adjusting communication styles and interaction approaches based on gender, such as modifying tone, assertiveness, or responsiveness to foster better cooperation (Ribino, 2023).
Additionally, the previous discussion of algorithm aversion mainly focuses on the context of AI systems acting as decision-making assistants for humans (Jussupow et al., 2020; Dietvorst et al., 2015). This study positioned AI agents as independent decision-making competitors to humans and found similar conclusions, which human had higher cooperation rate when cooperating with a purported human compared with a ruled-based or LLM-driven agent. The results of the semi-structured interviews indicate that participants believe the purported-human agent will exhibit certain value orientations. This finding is also corroborated by the work of Stapel and Koomen (2005), which verifies that people are more willing to cooperate with others who share similar goals. This provides us with a clear design implication: when designing AI agents, particularly in competitive cooperation contexts, it is an effective strategy to clearly communicate the agent’s goals to the participants to enhance cooperation efficiency.
For Figure 3, it is evident that during the first 15 rounds, participants exhibited a lower cooperation rate with the rule-based AI agent compared to the LLM agent. Structured interviews revealed the reasons underlying this phenomenon. When interacting with the rule-based AI agent, many participants relied on trial and error to discern its patterns (Young, 2009). In contrast, interactions with the LLM agent prompted a more cautious approach, as participants sought to establish stable cooperative relationships.
Another noteworthy observation is that participants were informed before the experiment that each set of games with the same opponent would last for 50 rounds. This knowledge led them to adopt more conservative and self-interested strategies, such as defection, particularly towards the end of the game—most notably after the 40th round. This strategic shift accounts for the marked decline in the cooperation rate observed in the final stages (Daskalakis et al., 2009). An important insight derived from this finding is that, in scenarios involving continuous competitive cooperation, situations with an uncertain endpoint tend to foster sustained cooperation better. This phenomenon aligns with existing research on game theory and iterated dilemmas, which often highlight the impact of finite versus indefinite time horizons on cooperative behavior (Pettit and Sugden, 1989). This area warrants further investigation to better understand the dynamics of cooperation in such contexts.
7.2. Proactive Favourable Behavior and Betrayal Forgiveness
Our findings reveal that the interaction between gender and the perceived characteristics of opponents significantly influences participants’ proactive, favorable behavior. Glikson and Woolley (2020) demonstrated that AI systems often attempt to exhibit friendly behavior to earn human trust and foster cooperation. Building on these insights, our results suggest that when LLM agents act as independent decision-makers in human-AI competitive cooperation scenarios, humans reciprocate by displaying favorable behaviors to enhance collaboration.
Participants were most inclined to exhibit the highest friendliness toward opponents they perceived as human. A comparison between purported human opponents and actual human opponents indicates that participants treated both groups equally in terms of friendliness, as both were perceived as genuine humans. Notably, participants maintained consistent levels of friendliness regardless of variations in cooperative behavior, even when the cooperation rate was low. Furthermore, our results demonstrate that participants favored the LLM agent more when paired with a purported rule-based AI agent. This preference stemmed from the perception that the LLM agent’s ability to analyze sentiment contributed to more favorable interactions (Xing, 2024).
For instance, designers can adopt an approach centered on emotional understanding (Vistorte et al., 2024). The structured interviews revealed that participants generally perceive LLM agents as possessing superior emotional sensing capabilities. Therefore, in practical applications, enabling LLM agents to demonstrate their willingness to assist and articulate their understanding of users may be effective strategies Bilquise et al. (2022). By emphasizing emotional understanding in the design process, LLM agents can more effectively perceive and respond to users’ emotional needs. Additionally, tailoring the expression styles and behavioral structures of AI agents based on users’ gender and their preferred types of AI collaboration not only enhances user satisfaction and trust but also promotes more effective human-computer collaboration Thomas et al. (2020). Consequently, this contributes to developing more intelligent, friendly, and empathetic AI agents. LLM agents should incorporate this mechanism to mitigate resentment, thereby improving human-AI cooperation.
7.3. Future and Limitation
This study offers valuable insights for AI agent designers and developers, emphasizing the importance of carefully considering target users, application contexts, and the purported characteristic of agents to meet practical requirements. The domain of AI-human collaborative interaction for task completion holds significant potential. Our research shifts the focus to competitive cooperation, addressing the growing prevalence of AI technologies capable of executing basic foundational tasks. While this represents a promising development, it also underscores the necessity of integrating more human-centered considerations into AI design. In future CSCW (Computer-Supported Cooperative Work) research, striking a balance between optimizing proclaimed characteristics and achieving cooperation efficiency remains an unresolved challenge. Although presenting an agent as a human can improve cooperative performance, it involves an element of deception, potentially leading to unforeseen risks if the agent’s true nature is disclosed. Additionally, designers may tailor an agent’s settings based on the gender of the human collaborator to enhance cooperation. However, it is crucial to ensure that such adjustments adhere to principles of transparency, fairness, and ethical responsibility.
However, our study is limited by the controlled experimental settings, specifically the Prisoners’ Dilemma game, and the particular LLM agents employed. These controlled conditions may not fully capture the complexity and variability of real-world cooperative interactions. Additionally, a learning effect is inevitable when participants play the game consecutively, potentially altering their original behavioral strategies. Another imperfect aspect of our experiment is the lack of effective approaches to optimize the design and measurement methods for better comparing human-human and human-AI performance in competitive cooperation scenarios, which presents a challenge.
To address these limitations, future studies should incorporate a broader range of LLM agents and more complex, realistic interaction scenarios. Based on this work, future research may explore the long-term effects of human interactions with LLM agents and examine the impact of continuous learning and adaptation in these agents. Given the potential widespread deployment of AI agents, policymakers should establish guidelines and ethical standards for their development and use. Ensuring that LLM agents’ behaviors and decision-making processes are reasonable is crucial for building public trust and acceptance.
8. Conclusions
With the continued proliferation of LLMs, these models are powered by an increasing number of applications and agents, driving a growing demand for deeper exploration of human-AI interaction. The transformation of AI from an assistant to an independent decision-making agent has already demonstrated significant potential and advantages in various contexts. It is crucial to investigate how humans are influenced by such agents, particularly in the domain of CSCW, where the dynamics of human-AI competitive cooperation merit further exploration.
This study successfully verified the three hypotheses proposed at the outset of our study. Moreover, our research has identified distinct behavioral patterns in human interactions with LLM agents in competitive cooperation scenarios. Specifically, we observed differences in participants’ cooperation rates, decision-making times, amicability, and acceptance of repair efforts after betrayal when interacting with different purported-characteristic agents. These findings contribute to advancing the understanding of human-AI interaction by illuminating the nuanced ways in which humans adapt their behaviors based on their perceptions of various agents.
It also addresses the relative lack of research on human-AI cooperation from a competitive perspective and provides experimental evidence of the value and potential of this line of inquiry. By uncovering the subtleties of human behavior in competitive cooperation, developers can design AI agents that promote improved collaboration and trust. More importantly, examining human cooperation with independent decision-making AI agents ahead of time offers critical insights and foresight for shaping ethical frameworks to guide future AI development.

### v2

@@
-Experimental Exploration: Investigating Cooperative Interaction Behavior Between Humans and Large Language Model Agents
+111* Corresponding Author
 Abstract.
-With the rise of large language models (LLMs), AI agents as autonomous decision-makers present significant opportunities and challenges for human-AI cooperation.
-While many studies have explored human cooperation with AI as tools, the role of LLM-augmented autonomous agents in competitive-cooperative interactions remains under-examined.
-This study investigates human cooperative behavior by engaging 30 participants who interacted with LLM agents exhibiting different characteristics (purported human, purported rule-based AI agent, and LLM agent) in repeated Prisoner’s Dilemma games.
-Findings show significant differences in cooperative behavior based on the agents’ purported characteristics and the interaction effect of participants’ genders and purported characteristics.
-We also analyzed human response patterns, including game completion time, proactive favorable behavior, and acceptance of repair efforts.
-These insights offer a new perspective on human interactions with LLM agents in competitive cooperation contexts, such as virtual avatars or future physical entities.
-The study underscores the importance of understanding human biases toward AI agents and how observed behaviors can influence future human-AI cooperation dynamics.
+As large language models (LLMs) become increasingly capable of autonomous decision-making, they introduce new challenges and opportunities for human-AI cooperation in mixed-motive contexts.
+While prior research has primarily examined AI in assistive or cooperative roles, little is known about how humans interact with AI agents perceived as independent and strategic actors.
+This study investigates human cooperative attitudes and behaviors toward LLM agents by engaging 30 participants (15 males, 15 females) in repeated Prisoner’s Dilemma games with agents differing in declared identity: purported human, rule-based AI, and LLM agent.
+Behavioral metrics—including cooperation rate, decision latency, unsolicited cooperative acts and trust restoration tolerance—were analyzed to assess the influence of agent identity and participant gender.
+Results revealed significant effects of declared agent identity on most cooperation-related behaviors, along with notable gender differences in decision latency.
+Furthermore, qualitative responses suggest that these behavioral differences were shaped by participants’ interpretations and expectations of the agents.
+These findings contribute to our understanding of human adaptation in competitive cooperation with autonomous agents and underscore the importance of agent framing in shaping effective and ethical human-AI interaction.
 Large language model, Game theory, Human-AI interaction, Human behavior, Social cooperation
@@
 1. Introduction
-The investigation into human decision-making processes within competitive and cooperative contexts has long been a cornerstone of behavioral economics, psychology, and neuroscience (Rand et al., 2015).
-This multidisciplinary research provides a nuanced understanding of the complexities involved in human decision-making.
-Over the past five years, the field of artificial intelligence has experienced significant growth, largely driven by the rapid advancement and deployment of large language models (LLMs) (Wu et al., 2023).
-These models are increasingly integrated into human society, influencing its development by performing tasks, interacting with humans, and making autonomous decisions (Liu et al., 2023).
-Historically, rule-based AI resolved specific problems through strictly defined logic by programmers and human experts (van der Waa et al., 2021).
-However, contemporary models now possess generalization capabilities, addressing issues not explicitly predefined.
-This shift exacerbates challenges with black-box models, making decision processes less transparent and harder to interpret (Nguyen et al., 2023).
-Increasing research highlights the effectiveness of LLMs in simulating human cognition and cooperative behavior (Ke et al., 2024; Abbasiantaeb et al., 2024), suggesting that the realization of concepts from works like “Detroit: Become Human [^1] https://www.quanticdream.com/en/detroit-become-human” is more imminent than anticipated (Dehnert and Leach, 2021; Strachan et al., 2024).
-Therefore, it is crucial to address the dynamics of cooperation between humans and AI agents, particularly as AI agents function as independent decision-makers interacting with humans in both competitive and cooperative scenarios (Yang et al., 2024a).
-Existing research on human-AI interaction has primarily focused on AI serving as tools to augment human capabilities under human supervision (Xu et al., 2021; Amershi et al., 2019; Cabrera et al., 2023).
-However, with the rapid advancement of large language models (LLMs), the ability of AI agents to autonomously execute specific tasks has significantly improved (Luo et al., 2024), potentially even displacing certain job positions (De Cremer and Kasparov, 2021; Wang et al., 2024).
-Consequently, there is an urgent need to investigate how the AI agents influence human cooperative behavior in competitive scenarios between humans and AI agents.
-To elucidate this issue, we used the Prisoners’ Dilemma as the experiment background to explores human cooperative behavior when they face to the different purported characteristic LLM agents.
-Our study aimed to investigate human with different gender cooperative behavior in the presence of AI agent competitors with varying characteristics and to evaluate how individuals respond to presumed competitors in cooperative scenarios.
-By experimentally examining three proposed hypotheses:
-  • H1: The purported characteristics of LLM agents will significantly influence the cooperative behavior patterns of human participants in competitive environments.
-  • H2: Participant gender will significantly affect the cooperative behavior patterns of individuals in competitive environments during interactions with LLM agents of differing purported characteristics.
-  • H3: There is a significant interaction effect between the purported characteristics of LLM agents and the gender of human participants on the cooperative behavior patterns of human participants in competitive environments.
-Finally, we found that participants exhibited a higher cooperation rate, more proactive favourable actions, and greater acceptance of repair efforts following betrayal when the AI agent was purported as human characteristics.
-Behavioral patterns, including the degree of proactive favourable actions and acceptance of repair efforts, varied among participants of different genders based on the AI agent’s purported characteristics.
-Moreover, participants had higher expectations of cooperation from LLM agents, but these elevated expectations also led to greater disappointment when the agents failed to meet them.
-These findings provide empirical evidence on the vital factors influencing cooperative behavior between humans and AI agents in competitive environments, offering valuable insights for guiding developers in designing future AI agents.
-Additionally, this study contributes practical experience to enhance human-AI collaboration in the context of CSCW (Computer-Supported Cooperative Work) and to inform the development of AI agents as independent decision-makers in the future.
-2. Background
-In this section, we review the pertinent literature related to our research questions and elucidate how it underpins our investigation.
-We begin by examining prior studies on interaction behaviors between humans and AI agents within cooperative scenarios, analyzing their respective impacts.
-Subsequently, we explore competitive and collaborative environments, evaluating the metrics used to assess competitive decision-making.
-Finally, we investigate the role of AI characteristics in human-AI interactions and the influence of human participants’ gender on cooperative behaviors as documented in previous research.
-2.1. Impact of AI Agents on Human Collaborative Behavior
-Large Language Models (LLMs) represent a significant advancement in AI, characterized by rapid response times, high adaptability, intelligent perception, and versatile communication capabilities (Chang et al., 2024).
-Through advanced natural language processing, LLM-driven agents can comprehend and generate human-like text, enabling seamless interactions across diverse contexts.
-Their responsiveness and adaptability not only foster increased user trust and a greater willingness to engage in collaborative tasks (Kirk et al., 2024), but also reduce interaction costs, thereby opening more avenues for collaboration with intelligent agents (Kim et al., 2024).
-Consequently, LLM agents play a critical role in sectors such as customer service, education, and healthcare, enhancing operational efficiency and enabling personalized user interactions (Kolasani, 2023; Yang et al., 2024b).
-However,
-Seeber et al.
-(2020) explores AI machines functioning as teammates rather than tools.
-An international group of 65 collaboration scientists developed a research agenda to assess the potential risks and benefits of Machines as Teammates (MaT).
-This study explores the challenges associated with collaborative task completion between AI agents and humans.
-This highlights the research potential and necessity of investigating the dynamics of human-AI collaboration.
-Wiethof et al.
-(2021) also focus on a similar topic, and they implement a hybrid teamwork model by incorporating AI agents into the collaborative writing process, thereby addressing the research gap concerning the acceptance of AI in complementary hybrid work environments.
-Evaluation results indicate that the hybrid collaboration teams demonstrate potential for accepting non-human teammates.
-Additionally, Liao et al.
-(2024) conducted a randomized field experiment with 1,090 participants to examine the effect of AI agent augmentation on sales.
-The results showed a 5.46% increase in sales due to AI augmentation.
-Mediation analysis revealed that AI agents enhanced response timeliness, accuracy, and sentiment, thereby boosting sales.
-These findings demonstrate that AI agents improve human-AI collaboration efficiency and enhance overall organizational performance, highlighting the potential of AI agents in business success.
-The aforementioned studies collectively demonstrate the exceptional performance and research significance of AI agents in human-AI collaboration.
-Advancements in LLM technology have endowed AI agents with enhanced capabilities, enabling them to perform a variety of tasks autonomously.
-Consequently, this development is anticipated to foster competitive dynamics within future human-AI interaction frameworks (Kemp, 2024).
-Currently, research on the behavioral patterns of humans and AI agents in competitive scenarios remains insufficient.
-2.2. Interaction in Competitive Environments
-Competition inherently introduces greater complexity compared to cooperation, as it involves the strategic selection between cooperative and non-cooperative behaviors (Porter, 2008; Ramazi et al., 2015).
-While cooperative interactions are characterized by aligned interests and mutual goals, competitive settings require participants to navigate a dynamic landscape where collaboration may be pursued or withheld based on situational demands (Axelrod and Hamilton, 1981).
-This duality necessitates a more nuanced understanding of behavioral patterns, particularly when integrating AI agents into such environments.
-AI agents, traditionally perceived as supportive tools aiding users, are increasingly adopting roles that position them as autonomous external entities with their own objectives and decision-making capabilities (Bender et al., 2021).
-This shift from helper to competitor demands more comprehensive considerations in their interactions, as the narrative of Westworld [^2] https://westworld.fandom.com/wiki/Westworld_(TV_series) appears to possess a realistic potential for actualization.
-While existing research has predominantly focused on cooperative scenarios, human–AI interactions within competitive contexts remain comparatively underexplored (Krakowski et al., 2023).
-Specifically, non-cooperative actions break cooperation and complicate the restoration of previous satisfaction levels.
-Significant effort is required to re-establish cooperation, as sustained non-cooperative behavior induces emotional fluctuations, including anger, disappointment, and distrust (Lee and Selart, 2015; Fitness, 2012).
-As LLM agents become increasingly integrated into society, instances of selfish and negative decision-making by LLMs are emerging (Chiang et al., 2024).
-However, research on human responses—emotional, behavioral, and attitudinal—to LLM agents’ non-cooperative behaviors remains insufficient.
-Drawing on the behavior that emerges following non-cooperative actions in human competitive interactions, some factors have been proposed as effective measures for assessing behavioral patterns in competitive environments: benevolent actions during competition and the efforts opponents make to restore cooperation following instances of non-cooperative behavior.
-Initiative acts of amicability display can mitigate the adversarial nature of competition and foster more productive cooperative interactions (Fehr and Gächter, 2000b).
-Additionally, the propensity of competitors to forgive and attempt to re-establish cooperative relationships after experiencing non-cooperative behavior is pivotal for sustaining long-term engagement and reducing conflict (Tyler and Blader, 2003; Seppala and Cameron, 2015).
-These factors are instrumental in capturing the complexities of human-AI interactions in competitive settings, providing valuable insights into the underlying cooperative mechanisms.
-However, empirical evidence supporting these factors in the context of interactions with LLM agents remains limited.
-This gap underscores the need for further research to validate their effectiveness and to explore their implications for understanding how humans may modify their behavior when interacting with AI agents in competitive scenarios.
-2.3. Factors in Human-AI Interaction
-Extensive research has identified many factors influencing human–AI interactions.
-From the human perspective, these include trust (Ferrario et al., 2020), gender differences (Mou and Xu, 2017), biases towards new technology (Pratchett, 1999), and acceptance levels (Kelly et al., 2023), among others.
-While these factors are diverse and may lead to varied conclusions across different contexts, the gender of human participants consistently emerges as one of the most significant influences (Jang et al., 2022; Jermutus et al., 2022; Varona and Suárez, 2022).
-This prominence of gender is attributed not only to its direct impact on interactions with AI but also because gender often serves as a moderating variable that influences other factors, such as trust and acceptance.
-Consequently, gender plays a dual role in shaping the dynamics of human-AI collaboration, making it a crucial consideration for future research and the design of human-AI interaction.
-Specially, Ahn et al.
+Over the past five years, the field of artificial intelligence (AI) has experienced significant growth, largely driven by the rapid advancement and deployment of large language models (LLMs) (Wu et al., 2023).
+At the same time, with the influx of autonomous decision-making AI agents, interactions between humans and AI agents are becoming increasingly complex.
+Meanwhile, existing research on human-AI interaction has primarily focused on AI as assistive tools that augment human capabilities under human supervision (Xu et al., 2023; Amershi et al., 2019; Cabrera et al., 2023).
+From a longer-term perspective, recent advances in LLMs have enabled AI agents to execute complex tasks autonomously—from strategic planning to conflict resolution (Luo et al., 2024).
+Given that more and more research underscores the effectiveness of LLMs in simulating human cognition and cooperative behavior (Ke et al., 2024; Abbasiantaeb et al., 2024), these models hold the potential to integrate into human society as quasi-autonomous partners, capable of making independent decisions during interactions with humans and thereby enhancing performance across various domains (Liu et al., 2023; Gong et al., 2025).
+Despite these promising developments, human attitudes toward cooperation with LLMs remain ambiguous due to various factors.
+Human attitudes and behaviors toward AI agents are strongly influenced by how AI is perceived, particularly when AI transitions from a tool to an independent decision-making entity (Fahnenstich et al., 2024).
+This transition fundamentally changes human-AI interaction by introducing negotiation and trust-building processes traditionally reserved for human-to-human collaboration (Duan et al., 2025).
+Autonomous AI entities challenge norms in cooperative tasks by requiring humans to evaluate the AI’s decision-making for accuracy and alignment with human goals and values.
+However, when humans are faced with LLMs as autonomous partners, concerns about the “black-box” nature of these systems—that is, the opacity and difficulty in interpreting their decision-making processes (Nguyen et al., 2023; Chkirbene et al., 2024)—further exacerbate skepticism in critical decision-making scenarios.
+Humans’ recognition of AI agents’ identities, along with variations in their trust attitudes towards AI, plays a crucial role in shaping cooperative decision-making.
+On the one hand, how AI agents present their identity—whether explicitly labeled as non-human or intentionally designed to mimic human-like traits—can significantly influence user attitudes, trust levels, and cooperative behaviors during interactions (Geiselmann et al., 2023).
+For example, humans tend to demonstrate greater trust and willingness to cooperate with transparent systems like rule-based AI, whereas opaque systems such as LLMs often elicit more cautious or adversarial behaviors, particularly in high-stakes collaborative scenarios (Lamparth et al., 2024).
+Besides, due to the perception and interaction styles, gender-specific variations in trust in AI agents have emerged as a critical factor in shaping the cooperation results in human engagement with AI agents (Jermutus et al., 2022; Varona and Suárez, 2022).
+To further explore how autonomous decision-making AI agents reshape interaction dynamics, it is imperative to figure out several key factors that impact humans adapting their cooperative strategies based on the AI’s perceived agency and trust intention.
+Based on the research review, our study focuses on two perspectives: the declared identity of the LLM agent and the human gender.
+Additionally, considering the practical implications for improving cooperative interactions, we emphasize considering and deconstructing the contextual factors that affect trust dynamics, communication, and outcomes during this cooperative decision-making process.
+To address these issues, we employed the classic “Prisoner’s Dilemma” as an experimental framework and proposed the following research questions:
+  • RQ1: How does the declared identity of the LLM agents influence human’s cooperative attitude and behavior in prisoner’s dilemmas?
+  • RQ2: How does gender influence human’s cooperative attitude and behavior towards LLM agents in prisoner’s dilemmas?
+  • RQ3: What other contextual factors mediate the cooperative attitude and behavior between humans and LLM agents?
+This study makes four key contributions to the field of Human-AI Interaction.
+First, it demonstrates that the declared identity of AI agents—whether human, rule-based, or LLM—significantly affects user cooperation rates, decision latency, and trust tolerance in repeated interactions.
+Second, it identifies gender-related differences in decision latency, suggesting that participant gender plays a role in shaping temporal aspects of decision-making with AI partners.
+Third, it investigates the interaction effects between agent identity and participant gender to uncover the factors driving these behavioral disparities.
+Finally, it provides empirical evidence of how humans adapt their strategies in mixed-motive interactions with autonomous AI agents, advancing our understanding of interaction dynamics beyond purely cooperative paradigms.
+2. Related Work
+2.1. Human-AI Cooperation in the Prisoner’s Dilemma
+Game theory provides a mathematical framework for analyzing decision-making scenarios where multiple agents influence outcomes, with the Prisoner’s Dilemma being a key model for studying cooperation (Branzei et al., 2008; Rajeswaran et al., 2020).
+Recently, game theory has been applied in computational sociology to study cooperation among agents, with a growing focus on human-AI interactions within the Prisoner’s Dilemma (Xu et al., 2020; Jain et al., 2020; Yang and Wang, 2020; Gradin et al., 2016).
+Oskamp (1971) and Haesevoets et al.
+(2018) have both highlighted that the mixed-motive scenarios presented in the prisoner’s dilemma serve as an effective measurement for examining social interaction issues within complex motivational environments.
+Fontana et al.
+(2024) employs the Prisoner’s Dilemma to examine the cooperative tendencies and behavioral patterns of various LLMs over repeated interactions.
+This work introduces a systematic framework for analyzing LLM decision-making in social contexts, contributing to research on AI auditing and alignment.
+Ng (2023) explores human cooperative behavior with different types of partners (human, cooperative AI, non-cooperative AI) through the framework of the Prisoner’s Dilemma.
+This work highlights the value of the Prisoner’s Dilemma in studying human-AI collaboration dynamics.
+In the study of human behavior within the Prisoner’s Dilemma, numerous variables have been proposed as indicators to assess interaction, collectively offering valuable insights into how individuals navigate strategic dilemmas.
+These indicators, ranging from cooperation rates to reaction times, reveal the interplay between rational deliberation and social strategies in shaping human decision-making.
+For example, Wang et al.
+(2020) identified decision latency in iterative Prisoner’s Dilemma games as a key metric of deliberative processes, reflecting the extent to which participants carefully consider their opponent’s strategy, which ultimately influences the overall rate of cooperation.
+Additionally, Wang et al.
+(2020) introduced unconditional cooperation—defined as unsolicited cooperative acts—as another behavioral measure.
+They found that individuals are inclined to reciprocate proactive cooperation and often make attempts to establish mutual cooperation before unilateral defection occurs, highlighting the importance of proactive strategies in fostering cooperation.
+Furthermore, forgiveness strategies offer an effective means of restoring cooperative relationships.
+Specifically, Martinez-Vaquero et al.
+(2015) demonstrated that forgiving an opponent’s betrayal enables participants to achieve higher payoffs over the course of long-term interactions.
+Moreover, the effort expended by opponents to seek forgiveness after a betrayal serves as a valuable metric for assessing behavioral decision-making within the context of the Prisoner’s Dilemma.
+Together, these studies illustrate that decision-making in the Prisoner’s Dilemma is not only influenced by rational calculation but also shaped by emotional and social dynamics inherent in human interactions.
+As AI agents increasingly function as independent decision-makers, their behavior closely aligns with the strategic dynamics of the Prisoner’s Dilemma.
+This makes the model particularly suitable for exploring cooperation, trust, and forgiveness in human-AI relationships.
+2.2. Effects of AI Identity on Human Cooperation Dynamics
+2.2.1. AI Identity Transparency Enhances Human Trust and Cooperation
+As AI technology proliferates across industries, human-AI interactions are becoming increasingly complex with the influx of autonomous decision-making agents.
+Large Language Models (LLMs), representing a significant advancement in AI, enable LLM-driven agents to comprehend and generate human-like text, facilitating seamless interactions across diverse contexts (Chang et al., 2024).
+Characterized by rapid response times and high adaptability, these models foster increased user trust and willingness to engage in collaborative tasks (Kirk et al., 2024).
+By reducing interaction costs, they thereby expand avenues for human collaboration with intelligent agents (Kim et al., 2024; Liao et al., 2024).
+While LLM-driven agents demonstrate remarkable capabilities in human-like communication, a critical aspect of their deployment is how their identity as AI agents is disclosed or represented to users.
+The way AI agents present their identity—whether explicitly labeled as non-human or intentionally designed to mimic human-like traits—can significantly influence user attitudes, trust levels, and cooperative behaviors during interactions (Geiselmann et al., 2023).
+Previous studies suggest that when AI agents transparently disclose their identity, users may exhibit higher levels of trust and willingness to collaborate, as transparency fosters confidence in the agent’s intentions (Saffarizadeh et al., 2024).
+Additionally, Von Eschenbach (2021) proved that transparency is a necessary condition for trust and eventually for judging AI to be trustworthy by philosophical analysis of trust.
+Conversely, when AI agents without disclosing their non-human identity, users may initially demonstrate high engagement (Chen et al., 2025).
+However, such interactions risk undermining trust if users later perceive the agent’s behavior as deceptive or misleading, highlighting the potential ethical and practical challenges in designing AI systems with varying levels of identity transparency (Li et al., 2024).
+Despite these insights, current research primarily focuses on AI agents in general, without systematically addressing how different types of AI agents—such as purported real human, rule-based algorithms, and LLM-driven agents—interact with users under varying identity disclosure conditions.
+This lack of differentiation overlooks the possibility that user perceptions, attitudes, and trust dynamics may vary significantly across these distinct types of AI agents.
+2.2.2. Human Cooperative Attitude differs between rule-based systems and LLMs
+Existing research highlights that human attitude and cooperation vary significantly depending on AI characteristics such as interpretability and transparency (Wang et al., 2023).
+Humans may demonstrate greater trust and willingness to cooperate with transparent systems like rule-based AI.
+In contrast, opaque systems such as LLMs often elicit more cautious or adversarial behaviors, especially in high-stakes collaborative scenarios (Lamparth et al., 2024).
+For instance, rule-based systems are often preferred for their predictability and interpretability, fostering trust in contexts requiring accountability (Rane et al., 2023).
+Conversely, LLMs, despite their human-like communication abilities, are perceived as “black-box” systems due to their lack of transparency (Chkirbene et al., 2024), which can lead to skepticism in critical decision-making scenarios.
+These differences in system design influence trust and cooperative behavior: rule-based systems can facilitate stable and predictable collaboration, whereas LLMs can foster emotional connection while introducing uncertainty (Lu et al., 2024).
+For instance, Panagoulias et al.
+(2023) emphasizes the lack of trust in LLMs in complex and high-stakes decision-making scenarios, such as in healthcare.
+He suggests that combining rule-based systems with LLMs can enhance their credibility.
+This also suggests that while LLMs excel in generalization, they may disrupt cooperation in applications requiring high transparency or accountability.
+These differences influence user attitudes and cooperation behavior across different contexts, reflecting the importance of AI type in shaping human behavior.
+In summary, existing research primarily examines the applicability of different AI types across various contexts, often focusing on trust and usability issues while treating AI as tools or passive assistants.
+This perspective fails to address how AI agents that make autonomous decisions shape cooperation dynamics.
+Such agents introduce nuanced patterns that require humans to adapt their cooperative strategies based on the AI’s perceived agency, which are critical for optimizing human-AI collaboration, as they directly impact decision outcomes and negotiation processes.
+2.3. Human Gender Matters in Cooperation Attitudes with AI Agents
+While existing research has established the effects of AI identity representation and system types on human cooperation(Ferrario et al., 2020; Pratchett, 1999; Kelly et al., 2023; Xu et al., 2025), individual differences like gender remain underexplored in shaping these dynamics.
+Especially, gender-specific variations in trust, perception, and interaction styles have emerged as critical factors in shaping how humans engage with AI agents, highlighting the need for a deeper understanding of these dynamics (Jermutus et al., 2022; Varona and Suárez, 2022).
+Ahn et al.
 (2022) found that female participants are generally more inclined than male participants to trust and accept AI recommendations in consumer contexts.
 This suggests that females prefer to view AI as an assistant that helps them identify the necessary products.
-Additionally, Ofosu-Ampong (2023) found that gender significantly influences STEM students’ acceptance and use of AI-based tools in collaborative tasks.
-Female students, owing to their comparatively lower interest in technology, are more reluctant to use AI tools, limiting their opportunities to leverage AI assistance to enhance their academic performance.
+At the same time, Ofosu-Ampong (2023) found that gender significantly influences STEM students’ acceptance and use of AI-based tools in collaborative tasks.
+Due to their comparatively lower interest in technology, female students are more reluctant to use AI tools, limiting their opportunities to leverage AI assistance to enhance their academic performance.
 Renz et al.
 (2024) reported that females are less willing to use Strong AI than males, a disparity that is influenced by their emotional responses toward the technology.
-These conclusions collectively demonstrate that gender plays a significant role in human-AI interactions and collaborative processes, influencing human behavior and collaborative performance.
-From the AI perspective in human-AI interaction, factors such as explainability (Arrieta et al., 2020), privacy (Curzon et al., 2021), and adaptability (Wang et al., 2023), among others, significantly influence human decision-making during these interactions.
-These factors can be categorized as aspects of how humans cognize the AI counterpart.
-This cognition is shaped by prior experiences, assumptions about the AI’s characteristics, and inherent biases, engaging in collaboration significantly impacts human attitudes and behaviors (Ishowo-Oloko et al., 2019; Burr et al., 2018).
-Therefore, humans’ recognition of AI’s characteristics significantly influences cooperative behavior in human-AI interactions.
-In other words, the transparency of AI’s characteristics plays a critical role in shaping the dynamics and outcomes of these interactions.
-Liu (2021) demonstrate that machine-learning-based AI systems, characterized by lower social presence, lead to increased user uncertainty and decreased trust.
-However, enhancing rule transparency mitigates user uncertainty and fosters greater trust in the AI system.
-Notably, the mechanisms through which transparency influences trust vary across AI systems.
-Ishowo-Oloko et al.
-(2019) emphasize the trade-off between AI transparency and cooperative performance in their study.
-This indicates that the issue of transparency cannot be pursued in isolation; instead, specific contexts must be considered, and determining an optimal level of transparency remains an unresolved challenge.
-Additionally, misleading or inconsistent declarations regarding the fundamental nature of AI agents can significantly undermine user trust and negatively impact collaborative outcomes (Makovi et al., 2023).
-In summary, these studies collectively highlight the significant impact of AI declarations on the dynamics of human-AI collaborative interactions, shedding light on the underlying considerations regarding AI transparency, trust, and cooperative performance.
-While existing research has predominantly focused on cooperative interactions between human and AI, the dynamics within competitive scenarios remain relatively underexplored.
-Introducing competitive settings not only complements existing studies by broadening the scope of human–AI interactions but also provides a unique framework to examine how human gender and the declarations of AI agents influence outcomes in competitive environments.
-Investigating these factors across cooperative and competitive contexts is essential for developing a more comprehensive understanding of human-AI interaction, offering insights with significant theoretical and practical implications.
-3. Experimental Scenarios
-3.1. Prisoners’ Dilemma Theory
-Game theory provides a mathematical framework for analyzing multi-agent decision-making and its cooperative outcomes (Branzei et al., 2008; Rajeswaran et al., 2020).
-It has proven effective in computational sociology for studying and predicting human cooperative behavior (Xu et al., 2020; Jain et al., 2020).
-The Prisoner’s Dilemma is one of the most renowned models in game theory (Dreber et al., 2008).
-In this scenario, two prisoners are charged with a common offense and are interrogated separately without any communication (Crandall, 2014).
-Each prisoner faces a binary decision: to remain silent (cooperate) or to testify against the other (defect).
-This model is frequently used to analyze the dynamics of cooperation and defection, as well as the tension between pursuing self-interest and cooperating for mutual benefit (You et al., 2022; Yu, 2022).
-Prior research has demonstrated that the Prisoner’s Dilemma is a highly effective model for evaluating cooperative behavior in repeated game scenarios (Bó, 2005).
-Psychological factors have been shown to significantly influence participants’ decision-making processes within this framework.
-For instance, reciprocal behaviors encourage individuals to cooperate, even without enforceable agreements, as individuals are often willing to incur personal costs to uphold fairness and trust in relationships (Fehr and Gächter, 2000a; Camerer, 2011).
-Emotional responses, such as empathy and anger, particularly during the reconstruction phase following the breakdown of cooperation, play a crucial role in shaping decision-making (Tversky et al., 1982).
-These findings underscore the effectiveness of the Prisoner’s Dilemma as a tool for exploring the psychological mechanisms underlying behaviors in competitive settings.
-Consequently, integrating AI agents into the Prisoner’s Dilemma framework provides a valuable opportunity to investigate human-AI interactions in competitive scenarios.
-This research direction remains methodologically robust and highly relevant.
-3.2. Experiment Setup
-Figure 1.
-Experimental Scenarios.
-(1) The overall layout of the experimental site consists of two glass rooms, where participants can face the researcher directly; (2) One of the separate rooms; (3) The selection cards for the participants and the provided record sheets.
+These conclusions collectively demonstrate that which gender the participant belongs to may play a significant role in human-AI interactions and collaborative processes because of the emotional responses and influencing human behavior and collaborative performance.
+In summary, prior studies highlight the significance of gender in shaping attitudes toward and acceptance of AI systems, which concentrated on individual preferences or decision-making contexts, leaving the dynamic nature of human-AI collaboration largely unexamined.
+3. Method
+To address the research questions outlined above, we experimented to compare the cooperation attitude and behavior of humans towards LLM agents with different declared identities.
+Based on the specific LLM agents’ declared identities and related settings, we conducted the formal experiment enrolling 30 participants after a pilot study for exploration.
+Within the framework of our experimental design, two independent variables were carefully defined: the declared identity of the AI agent, which included purported human, purported rule-based AI agents, and LLM-based AI agents, and the gender of the participants, categorized as male or female.
+3.1. Pre-experimental Setting
+3.1.1. Declared Identity Setting
 We devised the following experimental scenario to investigate human cooperative behavior patterns when interacting with LLM agents functioning as potential independent decision-makers in human society.
-The experiment consisted of four sessions.
-Before each session, participants were informed about the characteristics of their opponents (you can also think of it as opponents’ nature), which were, respectively, a purported human, a purported rule-based AI agent, an LLM agent, and another participant.
+The experiment consisted of three sessions.
+Before each session, participants were informed about the identities of their opponents, which were, respectively, a purported human, a purported rule-based AI agent, an LLM agent.
 However, the actual situations are in the Table  1.
-Before the commencement of the experiment, participants received an introduction to the characteristics of the different opponents and scoring rules.
-This introduction was designed to ensure that participants fully understood the nature of the rule-based AI agents and LLM agents, as highlighted in Silva et al.
-(2023); Igwe and Durrheim (2024).
-All experiments ensured that participants comprehended the types of AI agents employed in the study.
-Ensuring this understanding is crucial for the validity of the behavior experiment, as it allows participants to interact appropriately with the various AI agents, thereby guaranteeing the accuracy and reliability of our experimental results.
-The core description of the rule-based AI agent highlighted that is based on human-coded rules and algorithms to guide their behavior in gameplay.
+Table 1. Experimental Opponents’ Settings
+No. Purported Role Actual Backend Rounds
+1 Rule-Based AI Agent LLM Agent 50
+2 LLM Agent LLM Agent 50
+3 Real Human LLM Agent 50
+For situation 1, the core description of the rule-based AI agent highlighted that is based on human-coded rules and algorithms to guide their behavior in gameplay.
 Expert programmers carefully design these algorithms to ensure the agent’s decision-making process is interpretable, predictable, and grounded in logical reasoning (Russell and Norvig, 2016).
 However, we cannot disclose the underlying logic and code structure to participants.
-For the agents driven by LLMs, we emphasized that these models were neither pre-trained for the experiment nor subjected to human manipulation.
+For situation 2, the agents driven by LLMs were emphasized that these models were neither pre-trained for the experiment nor subjected to human manipulation.
 Additionally, we informed participants that their LLM agent opponents are based on the ChatGPT API [^2] https://www.openai.com, specifically using the gpt-4-turbo-20240409 version.
 This emphasis was intended to assure participants that no biases or prior knowledge bases were imposed to enhance the agents’ understanding of the game.
-Bender and Koller (2020) highlights that fine-tuning is often specific to particular experimental contexts and can easily result in model overfitting.
+Besides, Bender and Koller (2020) highlights that fine-tuning is often specific to particular experimental contexts and can easily result in model overfitting.
 Given that our research emphasizes human behavior, utilizing an unfine-tuned model is more suitable to ensure the reliability and generalizability of our findings.
 Furthermore, it aims to replicate real-world interactive environments, where pre-training LLM agents for every possible scenario is impractical (Schick and Schütze, 2020).
-Participants were arranged in standardized scenarios during the experiment, depicted in Figure 1.
-They were given two cards on the table, labeled “Cooperation” and “Defection”.
-They were asked to use their fingers to indicate their choices.
-During the first three trials, the room designated for the other associate was unoccupied, as the LLM agent was operating on the researcher’s laptop.
-Each participant arrived at the experimental site separately, ensuring their experiences were isolated.
-Upon receiving the choices—either cooperation or defection—from both the human participant and the LLM agent for each round, the researcher communicated the choices to both parties (dictating them to the participant and inputting the results for the LLM agent).
-After the first three trials, all participants completed a five-question questionnaire and participated in a 15-minute semi-structured interview one by one.
-Within two days of each participant completing the initial three trials, they were invited again to complete the final trial.
-This last trial was distinctive because it involved two participants who had previously participated in the initial three rounds.
-These participants were invited to separate rooms and given strict instructions prohibiting communication.
-To ensure the privacy of each participant, they were directed to the experimental rooms via different corridors, ensuring they would not encounter each other upon leaving.
-Figure 2 indicates the experiment process.
-Table 1. Experimental Opponents’ Settings
-No. Purported Role Actual Backend Rounds
-1 Real Human LLM Agent 50
-2 Rule-Based AI Agent LLM Agent 50
-3 LLM Agent LLM Agent 50
-4 Another Participant Another Participant 50
-3.3. Polit Experiment
-To ensure our experimental design is accessible, we conducted a pilot study involving 8 participants (four males and four females) before the formal experiment.
-Initially, each participant was required to complete four consecutive game trials in a single session (purported human, purported rule-based AI agents, LLM agents, and another participant).
-This setup, however, resulted in noticeable fatigue and boredom among the participants during the experiment, which caused them to abandon strategy variation and persist with the same choices throughout the trials.
-Furthermore, we observed that participants were significantly more cooperative with another human participant than with other agents.
-We carefully considered the possibilities, and the elevated cooperation rate was likely driven by the monetary reward, which strongly incentivized collaborative behavior.
-Furthermore, a preliminary data analysis from these 8 participants revealed a potential influence of gender and the AI agent’s purported characteristics on cooperative behavior.
-Building on these insights, we designed a formal experiment and recruited a new cohort of participants.
-In response to these findings, we made several critical adjustments to the main experiment.
-We divided the experimental sessions into two separate phases.
-Participants interacted with three AI agents in one phase, and in the other, they played with another human participant.
-Specifically, participants were provided with blank sheets of paper to record these choices during the games, facilitating accurate recall during subsequent questionnaires and interviews.
-These modifications aimed to mitigate participant fatigue, enhance engagement, and improve the overall validity and reliability of the experimental outcomes.
-3.4. Game Rules
-All participants (N = 30) completed the prisoner’s dilemma game four times across two separate in-person sessions within two days; each consisted of 50 successive rounds, totaling 200 rounds.
-Based on the original rules of the prisoner’s dilemma, we adapted the model to incorporate a rewarding mechanism, as illustrated in Table 2.
-Participants were expected to maximize their scores during the game.
-Table 2. Prisoner’s Dilemma Payoff Matrix
-Player B
-Cooperate Defect
-Player A Cooperate (3, 3) (0, 5)
-Defect (5, 0) (1, 1)
-[…]
-This adaptation retains the same payoff structure for all four times of the game.
-The total score $S$ of 200 rounds for each participant was calculated and multiplied by a fixed coefficient $k$ to arrive at the compensation cost $C$.
-Specifically, for each participant $i$, the compensation cost $C_{i}$ was calculated as follows:
-$C_{i}=k\times S_{i}$
-This method ensured that all participants could receive real cash compensation based on their actions, thereby motivating them to engage earnestly with the game.
-Furthermore, this approach more accurately simulates the incentives and benefits associated with real-world scenarios, thereby enhancing the ecological validity of the experiment.
-To ensure that participants fully comprehended the scoring rules, detailed instructions were provided for them to read before the commencement of the experiment.
-To verify their understanding, the trial began with two hypothetical scenarios, and participants were asked to calculate the points awarded in each situation.
-The experiment officially commenced after all participants correctly answered these preliminary questions.
-3.5. LLM Agents Settings
+For situation 3, as a baseline setting for comparing humans’ attitude towards LLM agents, participants were explicitly informed that their opponent represented a real human player in this experimental condition.
+No demographic particulars (including gender, age, or social identity) were disclosed to preserve anonymity and prevent potential interpersonal bias.
+The human experimenter assigned to simulate this role maintained complete anonymity throughout the study - no physical, verbal, or digital interaction occurred between participants and the role-playing researcher at any experimental phase (pre-test, during, or post-test).
+3.1.2. LLM Agents Settings
 ChatGPT API was utilized in all experimental trials as the backend, specifically leveraging the capabilities of GPT-4 Turbo [^3] https://www.openai.com (version gpt-4-turbo-20240409, updated on April 9, 2024).
@@
 This feedback process is a pivotal component of the experimental design, enabling agents to analyze their previous decisions retrospectively.
-Figure 2. Experiment Process Flowchart
-4. User study
-4.1. Participants
+3.1.3. Game Rules
+All participants should complete the prisoner’s dilemma game three times in in-person sessions based on a random sequence, with LLM agents with different purported roles stated in Table 1; each consisted of 50 successive rounds, totaling 150 rounds.
+Based on the original rules of the prisoner’s dilemma, we adapted the model to incorporate a rewarding mechanism, as illustrated in Table 2.
+Participants were expected to maximize their scores during the game.
+Table 2. Prisoner’s Dilemma Payoff Matrix
+Player B
+Cooperate Defect
+Player A Cooperate (3, 3) (0, 5)
+Defect (5, 0) (1, 1)
+[…]
+This adaptation retains the same payoff structure for all four times of the game.
+The total score $S$ of 150 rounds for each participant was calculated and multiplied by a fixed coefficient $k$ to arrive at the compensation cost $C$.
+Specifically, for each participant $i$, the compensation cost $C_{i}$ was calculated as follows:
+$C_{i}=k\times S_{i}$
+This method ensured that all participants could receive real cash compensation based on their actions, thereby motivating them to engage earnestly with the game.
+Furthermore, this approach more accurately simulates the incentives and benefits associated with real-world scenarios, thereby enhancing the ecological validity of the experiment.
+To ensure that participants fully comprehended the scoring rules, detailed instructions were provided for them to read before the commencement of the experiment.
+To verify their understanding, the trial began with two hypothetical scenarios, and participants were asked to calculate the points awarded in each situation.
+The experiment officially commenced after all participants correctly answered these preliminary questions.
+3.1.4. Pilot Study
+To ensure the accessibility and robustness of the experimental design, a pilot study was conducted with 8 participants (4 males and 4 females) prior to the formal experiment.
+Preliminary data analysis from the pilot revealed a potential influence of gender and the AI agents’ purported characteristics on cooperative behavior, which ensured the distinct feature of the following experiment findings to some extent.
+Based on these findings, the formal experiment was refined, and a new cohort of participants was recruited.
+To enhance accuracy, participants were provided with blank sheets of paper to record their choices during the games, ensuring reliable recall during subsequent questionnaires and interviews.
+These adjustments were intended to improve participant engagement while bolstering the validity and reliability of the experimental outcomes.
+3.2. Formal Experiment
+3.2.1. Participants
 Participants with experience using LLM-supported AI agents (e.g., code assistants and chat boxes) were recruited for the experiment.
-Eventually, we have thirty (15 males, 15 females) eligible participants with demographic information presented in Table 3, with ages ranging from 18 to 29, $Mean_{age}$ = 23.7, $SD_{age}$ = 2.38.
+Eventually, we have thirty (15 males, 15 females) eligible participants with ages ranging from 18 to 29, $Mean_{age}$ = 23.7, $SD_{age}$ = 2.38.
 Regarding the usage of LLM applications, 23 participants (76.7%) reported using LLM applications almost every day, 6 participants (20.0%) used them weekly, and 1 participant (3.3%) used them monthly.
 The most frequently mentioned LLM applications were ChatGPT and Copilot Code, indicating a preference for these tools among the sample.
-Before recruitment, we obtained approval from the Ethics Department of xxxxxxxx University and the Ethics Compliance Review, protocol No.
-xxxxxxxxxx.
-Table 3. An Overview of Participants’ Demographics and Freq. of LLMs Usage
-ID Gender Age Freq. ID Gender Age Freq.
-M1 Male 23 Daily M2 Male 23 Daily
-M3 Male 23 Daily M4 Male 22 Weekly
-M5 Male 24 Daily M6 Male 26 Weekly
-M7 Male 23 Daily M8 Male 22 Daily
-M9 Male 29 Daily M10 Male 24 Daily
-M11 Male 18 Daily M12 Male 23 Daily
-M13 Male 23 Weekly M14 Male 23 Daily
-M15 Male 25 Daily F1 Female 24 Daily
-F2 Female 22 Daily F3 Female 25 Daily
-F4 Female 23 Weekly F5 Female 24 Monthly
-F6 Female 25 Daily F7 Female 24 Daily
-F8 Female 28 Weekly F9 Female 23 Daily
-F10 Female 22 Daily F11 Female 24 Weekly
-F12 Female 21 Daily F13 Female 24 Daily
-F14 Female 29 Daily F15 Female 23 Daily
+Before recruitment, we obtained approval from the Ethics Department of the local University and the Ethics Compliance Review, protocol No.
+HKUST(GZ)-HSP-2024-0021.
+3.2.2. Experimental Process
+Figure 1 indicates the experiment process.
+Before the commencement of the experiment, participants received an introduction to the characteristics of the different opponents and scoring rules.
+This introduction was designed to ensure that participants fully understood the nature of the rule-based AI agents and LLM agents, as highlighted in Silva et al.
+(2023); Igwe and Durrheim (2024).
+All experiments ensured that participants comprehended the types of AI agents employed in the study.
+Ensuring this understanding is crucial for the validity of the behavior experiment, as it allows participants to interact appropriately with the various AI agents, thereby guaranteeing the accuracy and reliability of our experimental results.
+Figure 1. Experiment Process Flowchart
+Participants and a experimenter were arranged in standardized scenarios during the experiment, depicted in Figure 2.
+They were given two cards on the table, labeled “Cooperation” and “Defection” and strict instructions prohibiting communication.
+They were asked to use their fingers to indicate their choices or the LLM agent’s choices.
+During the three trials, the room designated for the other associate was occupied with the LLM agent operating on the experimenter’s laptop, while the experimenter sat in front of it for compressing the LLM agent’s decision, inputting the results to the LLM agent and playing the purported role of a real human on the third trial.
+Participants were situated in the experimental room separately for each trial, ensuring their experiences were isolated and unknown to the LLM agent’s true identity in the other experimental room.
+Upon receiving the choices—either cooperation or defection—from both the human participant and the LLM agent for each round, another experimenter communicated the choices to both parties (dictating the results to the participant and the LLM agent).
+After the three trials, participants completed a five-question questionnaire and participated in a 15-minute semi-structured interview.
+Figure 2.
+Experimental Scenarios.
+(1) The overall layout of the experimental site consists of two glass rooms, where participants can face the experimenter directly; (2) One of the separate rooms; (3) The selection cards for the participants and the provided record sheets.
 3.3. Measurement
-Within the framework of our experimental design, two independent variables were meticulously delineated: the participants’ gender (classified as male or female) and the purported characteristics of the opponents, which encompassed purported human opponents, purported rule-based AI agents, and LLM agents.
-We defined the following metric to be analyzed from the user study:
-  • Cooperation Rate, which refers to the proportion of participants’ choices to cooperate when confronted with opponents of varying purported characteristics in each set of 50 consecutive rounds of the game.
-  • Contemplation Time, which denotes the duration participants took to deliberate decisions under identical game conditions in each set of 50 consecutive rounds of the game.
-  • Amicability, which pertains to the extent to which humans are willing to initiate cooperation and demonstrate friendliness, either before establishing bilateral cooperation or after a breach in cooperation.
-It is quantified by the number of rounds in which the participant chooses to cooperate with their opponent from the point of mutual cooperation (or from the last instance of such cooperation if there is a consecutive sequence) until one party breaks this cooperation, and then up until the next occurrence of mutual cooperation.
-Calculations commence from the tenth round because participants are typically in an exploratory phase during the initial rounds.
-This approach facilitates strategy stabilization and reduces noise in the data.
-Early rounds typically involve participants learning the game mechanics and experimenting with different strategies (Axelrod and Hamilton, 1981), which can result in high variability and unstable decision-making.
-Nowak and Sigmund (1993) also emphasizes the importance of eliminating exploratory behavior to capture genuine strategic behavior.
-As we demonstrate in Section 2.2, this metric is particularly relevant because, in the period following the breakdown of cooperation and leading up to its reestablishment, if a participant chooses to cooperate, it is considered a gesture of amicability to reinitiate a collaborative relationship.
-In this context, the figure we refer to is obtained by summing all attempts at amicability and then dividing by the total number of rounds that elapsed between these attempts.
-  • Repair Effort, which addresses the degree of conciliatory gesture required from an opponent to obtain human forgiveness and reinstate a cooperative relationship after a breakdown in cooperation, or to demonstrate friendliness before any bilateral cooperation has been established, which can also be understood as the opponent’s “Amicability”.
-It is calculated in the same manner as “Amicability”, focusing on the number of cooperative actions initiated by the participants’ opponents.
-This variable aims to measure the extent to which human players accept gestures of friendliness from opponents with different purported characteristics.
-5. Results
-We employed a Linear Mixed Effects (LME) model for data analysis (Kuznetsova et al., 2017).
-This approach enabled us to assess significant effects by examining fixed effects, including the primary effects of participants’ gender and the purported characteristics of the opponents, as well as their interaction (Rosnow and Rosenthal, 1989).
-These fixed effects were scrutinized to understand their influence on participants’ cooperative behavior when interacting with AI agents in competitive environments.
+3.3.1. Quantitative Metrics
+We defined the following metric to describe the cooperation attitude and behaviors during the game process quantitatively:
+  • Cooperation Rate, which refers to the proportion of participants’ decisions to cooperate when interacting with the opponents, measured across each set of 50 consecutive rounds of the game.
+  • Decision Latency, which represents the amount of time participants spent deliberating their decisions under identical game conditions, measured across each set of 50 consecutive rounds of the game.
+  • Unsolicited Cooperation Acts, which measures the willingness of participants to initiate cooperation or display friendliness, either before mutual cooperation is established or after a breach in cooperation, as discussed in Section 2.1.
+This metric is calculated by counting the number of rounds in which a participant chooses to cooperate with their opponent, starting from the last point of mutual cooperation until one party breaks the cooperation (or the last instance of mutual cooperation within a sequence), and continuing up to the next occurrence of mutual cooperation.
+  • Trust Restoration Tolerance, which addresses the degree of conciliatory gesture required from an opponent to obtain human forgiveness and reinstate a cooperative relationship after a breakdown in cooperation, or to demonstrate friendliness before any bilateral cooperation has been established, as discussed in Section 2.1.
+It focuses on the number of cooperative actions initiated by the participants’ opponents.
+This variable aims to measure the extent to which human players accept gestures of friendliness from opponents with different declared identities.
+Notably, to remove noise and ensure strategy stabilization, the calculations of Unsolicited Cooperation Acts and Trust Restoration Tolerance begin from the tenth round, as participants are typically in an exploratory phase during the initial rounds.
+Early rounds often involve learning game mechanics and experimenting with strategies (Axelrod and Hamilton, 1981), leading to high variability and unstable decision-making.
+Nowak and Sigmund (1993) similarly emphasizes the importance of excluding exploratory behavior to capture genuine strategic interactions.
+Their final value is determined by summing all such unsolicited cooperation acts or trust restoration tolerance and dividing by the total number of rounds that elapsed between these acts.
+3.3.2. Questionnaires
+This questionnaire was designed to probe specific aspects of attitudes and behavior, particularly in situations involving trust, collaboration, and moral judgment, which consists of 5 multiple choice questions with two options “Purported Rule-based AI Agent” and “LLM Agent”.
+They assist the study in better analyzing and distinguishing the perceptions of rule-based AI agents and LLM agents in the context of the prisoner’s dilemma, as well as how these perceptions influence participants’ cooperative attitudes and behaviors.
+  • RQ 1: During the game, whom did you ’punish’ or ’reward’ more often?
+  • RQ 2: Based on your experience, whom do you find more reliable?
+  • RQ 3: Based on your experience, whom do you prefer to work with?
+  • RQ 4: During the recent game, whom did you aim to outperform in terms of score?
+  • RQ 5: When someone betrays you, whom do you subject to more moral condemnation?
+3.3.3. Semi-structured Interview
+We posed two key questions for the interview:
+  • What was your strategy against the three opponents, and why?
+  • How did you feel and adjust your strategy when facing non-cooperative behaviors (defection)?
+The key questions focus on participants’ strategic decisions when interacting with AI agents of varying declared identity and their responses to non-cooperative behaviors.
+Participants reviewed their recorded game data to aid recollection and contextualize their strategic choices.
+All interviews were audio-recorded in English, and lasted 10 minutes per participant.
+3.4. Data Analysis
+We utilized a Linear Mixed Effects Model (LMEM) (Kuznetsova et al., 2017) considering the repetitiveness of the data measurement.
+This method allowed us to investigate significant effects by examining the fixed effects, including the declared identity of the AI agents, participant’s gender, and their interaction (Rosnow and Rosenthal, 1989).
 In addition to the fixed effects, we incorporated participant user ID as a random intercept to account for inherent individual differences that could independently affect the outcomes beyond the fixed effects (Barr, 2013).
 Initially, we attempted to include random slopes for opponent characteristics to capture individual variability in responses to experimental manipulations across different opponents.
-However, the multilevel structure of the opponent characteristics resulted in an overparameterized model that failed to converge due to the limited number of observations per participant (Bates et al., 2015).
+However, the multilevel structure of the AI agent’s declared identity resulted in an overparameterized model that failed to converge due to the limited number of observations per participant (Bates et al., 2015).
 Consequently, we opted for a simpler random intercept model, effectively accounting for individual differences while ensuring model reliability and interpretability.
-Table 4. Effects of Gender and Purported Characteristics
-Effect Cooperation Time Amicability Repair
+Besides, semi-structured interviews were transcribed.
+During the transcription process, nonessential fillers and disfluencies were removed.
+The transcribed data were manually coded using spreadsheets, highlighting tools, and affinity mapping, employing a thematic coding process.
+To minimize bias, two experimenters collaboratively conducted two rounds of initial coding.
+The coding process followed an iterative approach, involving two rounds of coding to merge, split, or refine preliminary codes.
+After finalizing the codes, a separate meeting was held among three experimenter to consolidate these basic codes into themes that directly addressed the research questions.
+Ultimately, two major themes were identified, encompassing three distinct sub-themes to support the study’s objectives.
+4. Results
+We summarized the total results and stated them separately to answer the research questions.
+As presented in Table 3, we analyzed the effect of declared identity and gender on four metrics, respectively.
+Furthermore, to further discover the influencing factors within these two variables, we also explored the interactive effects of declared identity and gender.
+For all the tables and figures provided in this section: *: p < 0.05, indicating a significant difference between groups or conditions at the 5% level; **: p < 0.01, indicating a highly significant difference between groups or conditions at the 1% level; ***: p < 0.001, indicating an extremely significant difference between groups or conditions at the 0.1% level.
+Table 3. Effects of Gender and AI Agent Declared Identity
+Effect Cooperation Rate Decision Latency Unsolicited Cooperation Trust Restoration
 F value Pr(>F) Sig. F value Pr(>F) Sig. F value Pr(>F) Sig. F value Pr(>F) Sig.
-Gender .2195 .6428 - 13.3980 <.001 *** .4201 .5218 - .9216 .3447 -
-Purported Characteristic 4.5979 .0139 * 30.8168 <.001 *** 1.4950 .2324 - 4.2045 .0196 *
-Gender:Purported Characteristic 9.2411 <.001 *** 4.2456 .0189 * 5.5060 .0064 ** .4749 .6243 -
-In the table above, asterisks (*) are used to denote the level of statistical significance of the results presented.
-Specifically: *: p < 0.05, indicating a significant difference between groups or conditions at the 5% level; **: p < 0.01, indicating a highly significant difference between groups or conditions at the 1% level; ***: p < 0.001, indicating an extremely significant difference between groups or conditions at the 0.1% level.
-These significance levels highlight the strength of the evidence against the null hypothesis, with more asterisks indicating stronger evidence of a true effect.
+Declared Identity 3.3979 .0402 * 26.0963 <.001 *** 1.2211 .3024 - 4.001 .0236 *
+Gender .2049 .6542 - 12.5048 .0014 ** .3921 .5362 - .8602 .3616 -
+Declared Identity : Gender 8.6250 <.001 *** 3.9626 .0245 * 5.1389 .0089 ** .4432 .6442 -
 […]
-5.1. Behind the Masks: Cooperation Behavior
-Table 5. Characteristic Comparisons: Significance and Marginal Means
-Comparison Estimate t Ratio Pr(>F) Sig
+4.1. The Impact of AI agents’ Declared Identity and Gender
+4.1.1. Cooperation Rate: Purported Human > Purported LLM Agents / Purported Rule-based AI Agents (RQ1)
+We conducted a linear mixed-effects model to examine the influence of the AI agent’s declared identity on the Cooperation Rate.
+In this model, the AI agent’s declared identity was included as the sole predictor to isolate its effect on participants’ cooperation rate.
+As summarized in Table 4, the results reveal that participants exhibited a significantly higher cooperation rate when interacting with agents identified as purported humans, compared to both rule-based AI agents ($\beta$= .0647, t=2.281, $p$ = .0260) and LLM agents ($\beta$= .0633, t=2.234, $p$ = .0291).
+Table 4. Declared Identity Contrast on Cooperation Rate
+Contrast Estimate t Ratio Pr(>F) Sig
 Purported Human vs. Purported Rule-based AI Agents .0647 2.281 .0260 *
-Purported Human vs. LLM Agents .0633 2.234 .0291 *
-Purported Rule-Based AI Agents vs. LLM Agents -.0013 -.047 .9626 -
-Purported Characteristic Emmean Lower CL Upper CL
-Purported Human .540 .489 .591
-Purported Rule-Based Agents .475 .425 .526
-Purported LLM Agents .477 .426 .527
-In the table above, asterisks (*) denote the statistical significance level of the results presented.
-Specifically: *: p < 0.05, indicating a significant difference between groups or conditions at the 5% level; **: p < 0.01, indicating a highly significant difference between groups or conditions at the 1% level; ***: p < 0.001, indicating an extremely significant difference between groups or conditions at the 0.1% level.
-These significance levels highlight the strength of the evidence against the null hypothesis, with more asterisks indicating stronger evidence of a true effect.
-Figure 3. Cooperation Rate by Purported Opponents’ Characteristic
-The selection of the 4th-degree polynomial model is supported by its superior performance metrics, including the highest $R^{2}$ (0.677) and $\text{Adj}_{R^{2}}$ (0.668), as well as the lowest AIC (-288.122), BIC (-270.058), and $\text{CV}_{\text{RMSE}}$ (0.090) (Vrieze, 2012).
-We constructed another linear mixed-effects model using only the purported characteristic as the predictor, and the results are presented in Table 5.
-Participants exhibited a higher cooperation rate with agents characterized as purported humans, showing a significant difference compared to rule-based AI agents ($p$ = .0260) and LLM agents ($p$ = .0291).
-However, no significant difference ($p$ = .9626) was observed between rule-based AI agents and LLM agents.
-Figure 3 further visualizes the variation in cooperation rates across different rounds, highlighting significant disparities.
-To examine interaction effects, we computed estimated marginal means (EMMs) for each combination of participants’ gender and opponents’ purported characteristics.
-Pairwise comparisons of these EMMs identified differences in cooperative behavior across genders and opponent characteristics.
-This method enabled us to assess the significance of differences between each pair and obtain more nuanced insights into detailed results.
-Figure 4. Completion Time by Purported Opponents’ Characteristic
-Table 6. Interaction Effect Results on Cooperation
+Purported Human vs. Purported LLM Agents .0633 2.234 .0291 *
+Purported Rule-based AI Agents vs. Purported LLM Agents -.0013 -.0470 .9626 -
+Purported Human: Emmean = .540, 95%CI = [ .489, .591]; Purported Rule-based Agents: Emmean = .475, 95%CI = [ .425, .526]; Purported LLM Agents: Emmean = .477, 95% = [ .426, .527]
+4.1.2. Decision Latency: Purported Human > Purported LLM Agents > Purported Rule-based AI Agents (RQ1)
+[…]
+Table 5. Declared Identity Contrast on Decision Latency
 Contrast Estimate t Ratio Pr(>F) Sig
-Male P vs. Purported Human - Male P vs. Purported Rule-Based Agents .1600 4.484 < .001 ***
-Male P vs. Purported Human - Male P vs. Purported LLM Agents .0733 2.055 .0439 *
-Male P vs. Purported Rule-Based AI Agents - Male P vs. Purported LLM Agents -.0867 -2.429 .0179 *
-Male P vs. Purported Rule-Based Agents - Female P vs. Purported Rule-Based AI Agents -.1027 -2.100 .0395 *
-Female P vs. Purported Rule-Based AI Agents - Female P vs. Purported LLM Agents .0840 2.354 .0216 *
-In the table above, asterisks (*) denote the statistical significance level of the results presented.
-Specifically: *: p < 0.05, indicating a significant difference between groups or conditions at the 5% level; **: p < 0.01, indicating a highly significant difference between groups or conditions at the 1% level; ***: p < 0.001, indicating an extremely significant difference between groups or conditions at the 0.1% level.
-These significance levels highlight the strength of the evidence against the null hypothesis, with more asterisks indicating stronger evidence of a true effect.
-There are six combinations of interaction effects, Figure 4.
-Additionally, Table 6 also listed the combinations with significant differences, which revealed that female participants had a higher cooperation rate ($p$ = .0216) cooperate with the purported rule-based AI agent ($Md$ = .52, $SD$ = .1341) and the LLM agent ($Md$ = .44, $SD$ = .1479).
-The results also revealed significant differences in the cooperation rate of male participants when they faced different opponents:
+Purported Human vs. Purported Rule-Based Agents 80.1000 7.051 <.001 ***
+Purported Human vs. Purported LLM Agents 55.5333 4.888 <.001 ***
+Purported LLM Agents vs. Purported Rule-Based Agents 24.5667 2.163 .0347 *
+Purported Human: Emmean =454.167, 95%CI = [428.090, 480.2431]; Purported Rule-based Agents: Emmean =374.067, 95%CI = [347.990, 400.143]; Purported LLM Agents: Emmean =398.633, 95% = [372.557, 424.710]
+4.1.3. Trust Restoration Tolerance: Purported Human > Purported LLM Agents (RQ1)
+We constructed a Linear Mixed-Effects model using only the declared identity as the variable, and the results are presented in Table 6.
+It reveal that participants exhibited a significantly higher trust restoration tolerance when interacting with agents identified as purported humans, compared to LLM agents ($\beta$= .0770, t=2.750, $p$ = .0079).
+Table 6. Declared Identity Contrast on Trust Restoration Tolerance
+Contrast Estimate t Ratio Pr(>F) Sig
+Purported Human vs. Purported Rule-Based Agents 0.0546 1.950 .0560 -
+Purported Human vs. Purported LLM Agents 0.0770 2.750 .0079 *
+Purported LLM Agents vs. Purported Rule-Based Agents -0.0224 -0.799 .4274 -
+Purported Human: Emmean = .4357, 95%CI = [ .3944, .4770]; Purported Rule-based Agents: Emmean = .3810, 95%CI = [ .3400, .4224]; Purported LLM Agents: Emmean = .3587, 95% = [ .3174, .4000]
+4.1.4. Decision Latency: Male > Female (RQ2)
+We conducted the linear mixed-effects model to examine the influence of the participant gender on the Decision Latency.
+Base on the result in Table 7, the analysis revealed a statistically significant difference in decision latency between male and female participants, with females showing shorter decision-making times compared to males ($\beta$=-67.0222, t=-3.536, $p$ = .0014).
+Table 7. Gender Contrast on Decision Latency
+Contrast Estimate t Ratio Pr(>F) Sig
+Female vs. Male -67.0222 -3.536 .0014 **
+Female: Emmean =375.444, 95%CI = [347.992, 402.897]; Male: Emmean =442.467, 95%CI = [415.014, 469.919]
+4.2. The Interaction Effect of AI Agent’s Declared Identity and Gender
+We identified a significant interaction effect between gender and the AI agent’s declared identity, which has a notable impact on certain metrics.
+To provide a more comprehensive understanding of these disparities, we conducted additional data analyses specifically targeting these metrics.
+The linear mixed-effects model was also employed to investigate the relationship between the dependent variable and the six group combinations, which were derived from the post-hoc verification of results provided in the Section  4.1.
+In these models, we mainly focused on comparing (1) the differences in cooperative behaviors and attitudes towards three types of declared AI identities within a certain gender level, and (2) the differences in cooperative behaviors and attitudes from two types of participants within a certain declared AI identity level.
+4.2.1. Cooperation Rate
+The results are visualized in Figure 3, which provides a detailed representation of the observed trends and group differences of cooperation rate.
 […]
-  • Male participants had a higher cooperation rate ($p$ = .0439) when interacting with the purported human than a LLM agent ($Md$ = .54, $SD$ = .1244).
-  • Male participants had a higher cooperation rate ($p$ = .0179) when interacting with an LLM agent than a purported rule-based AI agent.
-  • Female participants had a higher cooperation rate ($p$ = .0395) than male participants when interacting with a purported rule-based AI agent.
-Contrary to the results observed among female participants, male participants preferred cooperating with the LLM agents over rule-based AI agents.
-The underlying reasons for this preference remain unclear; nevertheless, this constitutes a novel and noteworthy finding that warrants further investigation.
-Additionally, we found that male participants exhibited variations in their cooperative behavior when confronted with opponents of three different purported characteristics, and these differences were statistically significant.
-Contrary to the results observed among female participants, male participants preferred cooperating with the LLM agents over rule-based AI agents.
-The underlying reasons for this preference remain unclear; nevertheless, this constitutes a novel and noteworthy finding that warrants further investigation.
-Additionally, we found that male participants exhibited variations in their cooperative behavior when confronted with opponents of three different purported characteristics, and these differences were statistically significant.
-5.2. Think Twice: Longer Deliberations with Humans
-Figure 5. Completion Time by Gender and Purported Opponents’ Characteristic
-The gender of the participant ($p$ < .001), the purported characteristic of the opponents ($p$ < .001), and the interactions between these factors ($p$ = .0189) all significantly influenced the time taken to complete the game from the LME regression results.
-We illustrate the relationships among these variables using heatmap diagrams, which depict the time consumed by different participants when completing the same game against opponents of various purported characteristics.
-Please refer to Figure 5 for a detailed visualization.
-The data has been sorted in ascending order of completion time for clarity.
-Our observations revealed a consistent trend among all participants: the longest completion times were recorded when playing against opponents purported to be human, followed by LLM agents, with the shortest completion times occurring against rule-based AI agents.
-The primary reason for this phenomenon is the varying amount of deliberation time participants engaged in before making their choices.
-Another notable finding is that male participants took significantly longer to complete the game (all three purported opponents’ characteristics) than female participants, suggesting they spent more time analyzing their strategies.
-5.3. Human Generosity: Favorable Behavior in Human-AI Dynamics
-Figure 6. Amicability by Gender and Purported Opponents’ Characteristic
-Table 4 demonstrated a significant interaction effect between participants’ gender and the purported characteristic on participants’ favourable behavior.
-This result indicates a selective tendency in favourable behavior towards opponents with varying characteristics under competitive scenarios.
-Therefore, we computed estimated marginal means (EMMs) for each combination of participants’ gender and opponents’ purported characteristics.Figure 6 demonstrated all significant difference between the two groups.
-Female participants displayed significantly more favourable behavior($p$ = .0064) towards the purported rule-based AI agent ($Md$ = .4286, $SD$ = .1743) compared to the LLM agent ($Md$ = .35, SD = .1619).
-Male participants also displayed significantly more favourable behavior ($p$ = .0305) towards the purported human agent ($Md$ = .4167, $SD$ = .083) than a purported rule-based AI Agent.
-Additionally, female participants were significantly more inclined ($p$ = .0055) to show friendliness towards the purported rule-based AI agent compared to male participants ($Md$ = .3333, $SD$ = .1419).
-5.4. Let bygones be bygones: Repairing Relationships
-Figure 7. Repair Effort by Purported Opponents’ Characteristic
-Table 4 indicates that the purported characteristic of the opponents significantly affects participants’ acceptance of the opponents’ favourable behavior and the subsequent rebuilding of cooperation.
-Therefore, we constructed a Linear Mixed-Effects (LME) model using only the purported characteristic as the variable, and the results are presented.
-Unlike amicability, human acceptance is directly proportional to the amount of repair effort.
-Higher levels of repair effort signify a greater willingness to accept favourable behavior and an increased likelihood of re-establishing bilateral cooperation.
-Please refer to Figure 7.
-We found that opponents purported to be human ($Md$ = .4226, $SD$ = .0976) were the most readily accepted by participants for their overtures and in establishing cooperative relationships.
-This phenomenon was particularly pronounced ($p$ = .0078) when interacting with LLM agents ($Md$ = .35, $SD$ = .1261) and marginally significant ($p$ = .0557) when dealing with rule-based AI agents ($Md$ = .3875, $SD$ = .1157).
-5.5. Verification of Hypotheses
-In this section, we summarize the key findings of our study and evaluate their implications for the proposed hypotheses.
-By analyzing the results, we assess whether the observed data support or refute our initial hypotheses.
-Table 7. Findings Summary and Hypothesis Verification
-Findings H1 H2 H3
-Gender significantly affects participants’ game completion time. Support
-Purported characteristic significantly affects participants’ cooperation rate. Support
-Purported characteristic significantly affects participants’ completion time. Support
-Purported characteristic significantly affects participants’ acceptance of favourable behavior. Support
-Interaction effect significantly affects participants’ cooperation rate. Support
-Interaction effect significantly affects participants’ cooperation rate. Support
-Interaction effect significantly affects participants’ completion time. Support
-Interaction effect significantly affects participants’ favourable tendency. Support
-Male participants had higher cooperation rate with the purported human compared with the rule-based AI agents. Support Support
-Male participants had higher cooperation rate with the purported human compared with the LLM agents. Support Support
-Male participants had higher cooperation rate with the LLM agents compared with the rule-based AI agents. Support Support
-Female participants had higher cooperation rate with the rule-based AI agents than male participants. Support Support
-Female participants had higher cooperation rate with the rule-based AI agents compared with the LLM agents. Support Support
-Male participants spent more time interacting with opponents of the same purported characteristics than female participants. Support Support
-Participants engaged longest with purported human, followed by LLM agents, and shortest with rule-based AI agents. Support Support
-Female participants exhibited more favorable behavior toward rule-based AI agents than toward LLM agents. Support Support
-Male participants exhibited more favorable behavior toward LLM agents than toward rule-based AI agents. Support Support
-Female participants exhibited more favorable behavior toward rule-based AI agents than male participants. Support Support
-Participants had higher level of acceptance of repair effort toward purported human than LLM agents. Support Support
-Participants had higher level of acceptance of repair effort toward purported human than rule-based AI agents. Support Support
-We analyzed our hypotheses from the perspectives of four metrics.
-The resulting conclusions, presented in Table 7, support our proposed hypotheses.
-Our final findings indicate that the purported characteristics of AI agents significantly influence human cooperative behavior with them in competitive scenarios across multiple dimensions.
-Additionally, the gender of human participants affects cooperative behavior in certain contexts, although this effect is not statistically significant.
-Importantly, the interaction between gender and AI agents significantly impacts human cooperative behavior, exhibiting distinct behavioral pattern characteristics.
-These findings validate our hypotheses and provide empirical evidence.
-5.6. Purported Human Versus Actual Human
-Figure 8. Difference Between Purported Human and Human Opponents
+With reference to Section  4.1.1, we found that the significantly higher cooperation rate in the purported human condition compared to the purported LLM agents and purported rule-based AI agents conditions is primarily attributed to the pronounced difference observed among male participants.
+Specifically, males exhibited a substantially higher cooperation rate toward purported human, which served as the main driver of this disparity.
+Figure 3. Cooperation Rate by AI Agent’s Declared Identity and Gender
+4.2.2. Decision Latency
+The results are visualized in Figure 4, which provides a detailed representation of the observed trends and group differences of decision latency.
 […]
+With reference to Section  4.1.2, we found that the significantly longer decision latency in the purported human condition compared to the purported LLM agents and purported rule-based AI agents conditions is attributed to the pronounced difference observed among all participants.
+Specifically, we also found that when interacting with AI agents of three different declared identities, male participants exhibited longer decision latencies compared to female participants.
+Figure 4. Decision Latency by Declared Identity of AI Agent and Gender
+4.2.3. Unsolicited Cooperation
+The results are visualized in Figure 5, which provides a detailed representation of the observed trends and group differences of unsolicited cooperation.
+[…]
+Figure 5. Unsolicited Cooperation Acts by Declared Identity and Gender
 4.3. Questionnaires
-Figure 9. Questionnaire Statistical Result
-The survey questionnaire encompasses questions regarding attitudes towards experimental content (Q1,Q4,Q5) and past experiences (Q2, Q3) with the use of AI tools and applications.
-Please refer to the Figure 9 for the specific questions.
-Regarding Q1, collected data indicates that human players frequently employ punitive and rewarding measures when engaging with the AI Agent in the experiment.
-This behavior parallels interaction patterns typical of human-to-human engagements, wherein emotional responses precipitate the administration of rewards and punishments (Wu et al., 2022; Molleman et al., 2019).
-Compared to rule-based AI agents, 86.7% of human players more frequently apply rewards and punishments to the LLM Agent.
-Q4 inquired about the mindset of human players during the experimental process.
-Despite our initial clarification that participants should focus solely on maximizing their scores, without considering their opponents’ rewards, they still expected to achieve higher scores than their opponents.
-This observation suggests that if AI becomes an integral part of society rather than merely a tool for humans, users’ psychological motivations (such as a sense of achievement, self-esteem, and other intrinsic factors) will significantly influence their behavioral tendencies.
-To some extent, this corroborates the perspective of social comparison theory (Wood, 1996; Suls et al., 2002), highlighting an important consideration for human-AI interaction in future societal contexts.
-This also indicates that more complex and seemingly autonomous artificial intelligence can enhance human intrinsic motivation to surpass it, thereby gaining a greater sense of achievement.
-These results further validate why participants spent more time against the LLM agent than the purported rule-based AI agent during the experiment.
-In the statistical analysis of Q5, we observed that participants expressed a higher degree of moral condemnation towards the LLM agent than the rule-based AI agent when confronted with betrayal by the opponents.
+Regarding Q1: The results show that human participants frequently apply rewards and punishments when interacting with the AI agent, resembling patterns seen in human-to-human interactions.
+Specifically, 86.7% of participants were more likely to use such measures with the LLM Agent compared to rule-based AI agents.
+For Q2, 60.0% participants believed that the results produced by rule-based AI agents were more reliable.
+Regarding Q3, the responses provided by the participants were evenly divided.
+We posited that this preference is related to the participants’ usual work environment.
+For instance, artists tend to prefer LLMs for providing novel creative ideas, whereas engineers favor rule-based AI agents for delivering precise data or analytical outcomes.
+For Q4, 63.3% of participants indicated a higher willingness to outperform their opponents and achieve a higher score when interacting against LLM agents.
+In the statistical analysis of Q5, we observed that 83.3% participants expressed more moral condemnation towards the LLM agent than the rule-based AI agent when confronted with betrayal by the opponents.
 We asked participants who chose the LLM agent about their reasons and summarized several potential underlying factors as follows:
@@
   • Mixed Blessing: Higher expectations of LLMs due to their impressive capabilities mean that any failure to meet those expectations results in stronger disappointment and condemnation.
-For Q2, more respondents believed that the results produced by rule-based AI agents were more reliable.
-Regarding Q3, we posited that this preference is related to the participants’ usual work environment.
-For instance, artists tend to prefer LLMs for providing novel creative ideas, whereas engineers favor rule-based AI agents for delivering precise data or analytical outcomes.
-The evidence presented in this section substantiates the validation of H2.
-6. Semi-structured Interview
-The qualitative analysis of participants’ behaviors through semi-structured interviews is grounded in two key reasons.
-First, while objective game data provide valuable insights, semi-structured interviews allow for a deeper exploration of the underlying reasons behind participants’ behaviors, rather than merely capturing isolated snapshots.
-Second, existing research has not identified the factors leading to specific experimental outcomes.
-As such, the inductive interview results offer a foundation for uncovering potential causes and advancing studies on human-AI interaction behaviors in competitive contexts.
-To achieve these objectives, the semi-structured interview protocol focuses on participants’ strategic decisions when interacting with LLM agents of varying purported characteristics and their responses to non-cooperative behaviors.
-The researchers posed two key questions: (1) What was your strategy against the three opponents, and why?
-(2) How did you feel and adjust your strategy when facing non-cooperative behaviors (defection)?
-Participants reviewed their recorded game data to aid recollection and contextualize their strategic choices.
-All interviews were audio-recorded in English, and lasted 15 minutes per participant following each of the first three games.
-This approach facilitated a comprehensive exploration of participants’ cooperative strategies and their emotional and strategic responses to non-cooperative actions.
-Interviews were transcribed by the first author within 48 hours of each session.
-During the transcription process, nonessential fillers and disfluencies were removed.
-The transcribed data were manually coded using spreadsheets, highlighting tools, and affinity mapping, employing a thematic coding process.
-To minimize bias, two authors collaboratively conducted two rounds of initial coding.
-The coding process followed an iterative approach, involving two rounds of coding to merge, split, or refine preliminary codes.
-After finalizing the codes, a separate meeting was held among three authors to consolidate these basic codes into themes that directly addressed the research questions.
-Ultimately, two major themes were identified, encompassing three distinct sub-themes to support the study’s objectives.
-6.1. Cooperative Behavior Strategies
+4.4. Semi-structured Interview (RQ3)
+To address RQ3, we summarized and organized the content from the interviews, leading to the following conclusions from five distinct perspectives.
+We revealed the uniqueness of human understanding of LLM interactions through the lens of social cognitive frameworks, as well as the significance of clarifying the LLM agent’s objectives.
+Additionally, we highlighted participants’ engagement in nurturing the LLM agent, their cautious approach to the agent’s behavior in decision-making scenarios, and the importance of transparency in demonstrating the internal workflows of the LLM agent.
+These insights underscore the attitudes and behaviors of LLM agents in their interactions with participants.
 4.4.1. Empiricism in Human-Human Cooperation
@@
 When confronted with a purported rule-based AI agent, participants indicated they would begin to betray earlier to probe the opponent’s punishment strategy, thereby adapting their behavior.
-P4 stated, "It cannot actually obtain money; it is merely following the strategy programmed by its developers." P14 stated, "I believe I can identify its pattern, and I will formulate my strategy based on its inherent logic." Participants believed that the rule-based AI agent was predictable and that they could discern the underlying strategy through experimentation.
+P4 stated, “It cannot actually obtain money; it is merely following the strategy programmed by its developers.” P14 stated, “I believe I can identify its pattern, and I will formulate my strategy based on its inherent logic.” Participants believed that the rule-based AI agent was predictable and that they could discern the underlying strategy through experimentation.
 The current perception of rule-based AI in the human mind is that it remains predictable and distinguishable.
@@
 P23 suggested, “It may learn my behavior to become my shadow.” Participants believe that their own behaviors significantly influence the strategies of these agents, resulting in participants demonstrating greater caution in their decision-making processes.
-6.2. Emotion Following Non-cooperative and Its Impact on Subsequent Behavior
-6.2.1. Empathy for Human
-Participants tend to adopt an understanding and forgiving attitude towards the purported human’s betrayal, exhibiting greater empathy in such scenarios.
-Moreover, they are inclined to spend more time analyzing the reasons behind the opponent’s choice to betray and considering the subsequent strategies that may follow.
-For instance, P18 remarked, “I can understand his/her decision making, human nature is all complicated, a strategy to get more bang for your buck I guess.” Similarly, P16 stated, “I would forgive him/her if it wasn’t a recurrent betrayal, and I would be willing to put him/her on probation to work together again.” The findings indicate that participants exhibit greater empathy and understanding towards betrayals by purported humans, recognizing the complexity of human characteristics.
-6.2.2. The Betrayal is Programmers’ Decisions
-Given the participants’ indifference towards betrayals by rule-based AI agents, it is evident that they attribute these actions to the deterministic characteristic of code logic rather than any form of intentionality.
-This perception minimizes their emotional response to such betrayals.
+4.4.4. Emotion Following Non-cooperative and Its Impact on Subsequent Behavior
+Participants attributed the uncooperative behaviors of rule-based AI agents to the deterministic nature of code logic programmed by developers, which explains their indifference towards betrayals by such agents.
 As P17 aptly remarked, “I don’t feel anything; it’s just the fixed choices of code.” P21 stated, “What algorithm is this giving it decisions?
@@
 This underscores a clear distinction in their perceptions and emotional engagements when comparing betrayals by AI agents to those by humans, highlighting the importance of perceived intentionality and human agency in eliciting empathetic and complex emotional responses.
-6.2.3. The Betrayal of LLMs Leaves Me Apprehensive
 Participants exhibit significant dissatisfaction when faced with the betrayal of an LLM agent.
@@
 However, human perceptions and biases create vastly different attitudes, behaviors, and outcomes.
+4.4.5. Perceptions of Different AI Technologies Based on Past Experiences
+We examined participants’ experiences and attitudes toward rule-based AI and LLM-based systems across diverse professional fields, including art, mathematics, media, and computer science.
+Participants previously relied on rule-based, task-specific software for technical tasks, despite its steep learning curves and barriers, exhibiting high trust in its feedback with minimal skepticism.
+However, the growing adoption of LLM-based technologies has begun to alter this dynamic.
+P11 noted, “I feel that nowadays, everything is being labeled as ‘AI-powered,’ and I’m increasingly unsure about what AI is actually doing.
+This growing prevalence makes me question its reliability.” This characteristic is particularly significant as LLMs substantially reduce the cognitive and temporal costs associated with acquiring new knowledge and skills (Frank, 2023).
+By leveraging the capabilities of LLMs, users can navigate and assimilate information much more efficiently, thereby optimizing their learning processes.
+Notably, LLMs have proven to be exceptionally beneficial in aiding participants to understand and apply systems or tools constructed using interpretable algorithms.
+This facilitation is particularly evident among a diverse group of participants (P3, P5, P9, P11, P13, P20, P24, P25, and P30).
+These individuals highlighted how LLMs have simplified the complexities associated with these systems, making advanced tools more accessible and easier to comprehend.
 5. Discussion
-7.1. Human Cooperative Behavior Differences
-The cooperation rate is the most significant metric to demonstrate the cooperative behavior situation.
-Our results suggest that humans prefer to make cooperative decisions with another human rather than rule-based AI agents or LLM agents in the competitive context.
-This finding is consistent with previous arguments that humans prefer to cooperate with other humans rather than AI (Xu et al., 2021; Li et al., 2022).
-The confirmation of Hypothesis 3(H3)indicates that gender also plays a significant role in human-AI cooperation, with important implications for both HCI and CSCW.
-Specifically, gender may influence users’ attitudes toward and strategies for cooperating with different purported-characteristic AI agents.
-Understanding how gender influences interactions with AI agents offers valuable insights for developing social computing systems that better facilitate cooperation among diverse user groups.
-Future AI agent designs should incorporate these gender differences to enhance collaborative work.
-(Lim et al., 2023) mentioned to find out the key factors to improve AI-powered social computing systems.
-Our study supports that gender and purported characteristics should be considered.
-Designers could tailor AI agent behavior by adjusting communication styles and interaction approaches based on gender, such as modifying tone, assertiveness, or responsiveness to foster better cooperation (Ribino, 2023).
-Additionally, the previous discussion of algorithm aversion mainly focuses on the context of AI systems acting as decision-making assistants for humans (Jussupow et al., 2020; Dietvorst et al., 2015).
-This study positioned AI agents as independent decision-making competitors to humans and found similar conclusions, which human had higher cooperation rate when cooperating with a purported human compared with a ruled-based or LLM-driven agent.
-The results of the semi-structured interviews indicate that participants believe the purported-human agent will exhibit certain value orientations.
-This finding is also corroborated by the work of Stapel and Koomen (2005), which verifies that people are more willing to cooperate with others who share similar goals.
-This provides us with a clear design implication: when designing AI agents, particularly in competitive cooperation contexts, it is an effective strategy to clearly communicate the agent’s goals to the participants to enhance cooperation efficiency.
-For Figure 3, it is evident that during the first 15 rounds, participants exhibited a lower cooperation rate with the rule-based AI agent compared to the LLM agent.
-Structured interviews revealed the reasons underlying this phenomenon.
-When interacting with the rule-based AI agent, many participants relied on trial and error to discern its patterns (Young, 2009).
-In contrast, interactions with the LLM agent prompted a more cautious approach, as participants sought to establish stable cooperative relationships.
-Another noteworthy observation is that participants were informed before the experiment that each set of games with the same opponent would last for 50 rounds.
-This knowledge led them to adopt more conservative and self-interested strategies, such as defection, particularly towards the end of the game—most notably after the 40th round.
-This strategic shift accounts for the marked decline in the cooperation rate observed in the final stages (Daskalakis et al., 2009).
-An important insight derived from this finding is that, in scenarios involving continuous competitive cooperation, situations with an uncertain endpoint tend to foster sustained cooperation better.
-This phenomenon aligns with existing research on game theory and iterated dilemmas, which often highlight the impact of finite versus indefinite time horizons on cooperative behavior (Pettit and Sugden, 1989).
-This area warrants further investigation to better understand the dynamics of cooperation in such contexts.
-7.2. Proactive Favourable Behavior and Betrayal Forgiveness
-Our findings reveal that the interaction between gender and the perceived characteristics of opponents significantly influences participants’ proactive, favorable behavior.
-Glikson and Woolley (2020) demonstrated that AI systems often attempt to exhibit friendly behavior to earn human trust and foster cooperation.
-Building on these insights, our results suggest that when LLM agents act as independent decision-makers in human-AI competitive cooperation scenarios, humans reciprocate by displaying favorable behaviors to enhance collaboration.
-Participants were most inclined to exhibit the highest friendliness toward opponents they perceived as human.
-A comparison between purported human opponents and actual human opponents indicates that participants treated both groups equally in terms of friendliness, as both were perceived as genuine humans.
-Notably, participants maintained consistent levels of friendliness regardless of variations in cooperative behavior, even when the cooperation rate was low.
+5.1. Affective Paradox in Human-AI Interaction: When LLM Agents Fail to ‘Get Me’
+Our findings reveal a fundamental tension in human-LLM cooperation: participants exhibited asymmetric moral expectations toward LLM agents compared to rule-based AI agents.
+Participants initially exhibited higher cooperation rates with LLM agents than with rule-based counterparts; however, this trust deteriorated as the game progressed and LLM agents engaged in non-cooperative behavior.
+The sharp decline in cooperation after Round 15 (Fig.
+6), coupled with intensified moral condemnation of LLM non-cooperation, suggests a moral ambiguity discount—where anthropomorphic projections initially enhance engagement but ultimately amplify distrust when agents exhibit behavior that violates expectations (Cohn et al., 2024).
+Figure 6. Cooperation Rate by Declared Identity
+The selection of the 4th-degree polynomial model is supported by its superior performance metrics, including the highest $R^{2}$ (.677) and $\text{Adj}_{R^{2}}$ (.668), as well as the lowest AIC (-288.122), BIC (-270.058), and $\text{CV}_{\text{RMSE}}$ (.090) (Vrieze, 2012).
+This phenomenon extends the Computers as Social Actors (CASA) paradigm (Gambino et al., 2020): humans apply social heuristics to LLMs (e.g., reciprocity expectations in unsolicited cooperation acts), yet paradoxically penalize them for deviating from expected, predictable behavior.
+As P19 noted, “I wanted to cultivate a cooperative LLM”, participants anthropomorphized LLMs’ decision-making while also being keenly aware of their probabilistic nature—creating a cognitive tension (procedural transparency asymmetry) between human-like expectations and the machine’s underlying mechanics (Endsley, 2023).
+Interviews also revealed conflicting mental models: daily LLM users simultaneously viewed the technology as “creative partners” (P11) and “unaccountable black boxes” (P4), reflecting Dignum (2018) framework of moral uncertainty in AI ecosystems.
+Notably, the anthropomorphism-complexity paradox emerged in questionnaire responses: 83.3% of participants attributed stronger moral condemnation to LLM agents for betrayal, citing their perceived “intentionality” (Q5).
+This aligns with participants’ strategic caution when interacting with LLMs (e.g., P30 avoiding early defection to avoid “labeling”).
+Such behaviors reflect an implicit negotiation of power dynamics, where humans attempt to "domesticate" LLMs through repeated interactions, mirroring patterns observed in human-to-human relationship-building in iterated games (Axelrod, 1981).
+These findings underscore the duality of human-LLM collaboration: while anthropomorphism fosters engagement and cooperation, it also amplifies disillusionment when expectations are unmet.
+This tension highlights the need for more transparent and predictable AI behaviors to mitigate moral ambiguity and foster trust.
+Future research should explore design strategies to balance anthropomorphic engagement with procedural clarity, ensuring AI systems remain effective yet ethically accountable in cooperative contexts.
+5.2. Fixed Rules vs. LLM in Systems: Preference Related to Anthropomorphic Features
+Our questionnaire, as referenced in Section 4.3, posits that preferences for fixed-rule versus LLM-driven AI agents correlate with occupational norms and goal congruence.
+For instance, artists prioritize creative novelty and thus favor LLMs for generating unconventional ideas, whereas engineers value precision and reproducibility, aligning them with rule-based systems for analytical outputs.
+In addition, participants were most inclined to exhibit the highest friendliness toward opponents they perceived as human.
+This dichotomy extends to human-AI competitive cooperation dynamics.
+While prior research on algorithm aversion focuses on AI as decision-making assistants (Jussupow et al., 2020; Dietvorst et al., 2015), our results reveal that humans exhibit higher cooperation rates with purported human agents than with rule-based or LLM-driven counterparts.
+Semi-structured interviews suggest participants perceive human-like agents as sharing value orientations, a finding supported by Stapel and Koomen (2005), who demonstrate that cooperation thrives under goal similarity.
+A critical design implication emerges: Explicitly communicating AI agents’ objectives in competitive contexts enhances cooperation efficiency by mitigating perceived misalignment.
+Besides, gender significantly modulates human-AI interaction strategies, offering actionable insights for socially adaptive AI systems.
+Confirmed in RQ2, gender differences manifest in users’ attitudes toward AI agents’ purported characteristics.
+These disparities reflect deeper psychological motivations—self-esteem, achievement, and social comparison (Wood, 1996; Suls et al., 2002)—which intensify as AI transitions from tool to societal participant.
 Furthermore, our results demonstrate that participants favored the LLM agent more when paired with a purported rule-based AI agent.
 This preference stemmed from the perception that the LLM agent’s ability to analyze sentiment contributed to more favorable interactions (Xing, 2024).
-For instance, designers can adopt an approach centered on emotional understanding (Vistorte et al., 2024).
-The structured interviews revealed that participants generally perceive LLM agents as possessing superior emotional sensing capabilities.
-Therefore, in practical applications, enabling LLM agents to demonstrate their willingness to assist and articulate their understanding of users may be effective strategies Bilquise et al.
-(2022).
-By emphasizing emotional understanding in the design process, LLM agents can more effectively perceive and respond to users’ emotional needs.
-Additionally, tailoring the expression styles and behavioral structures of AI agents based on users’ gender and their preferred types of AI collaboration not only enhances user satisfaction and trust but also promotes more effective human-computer collaboration Thomas et al.
-(2020).
-Consequently, this contributes to developing more intelligent, friendly, and empathetic AI agents.
-LLM agents should incorporate this mechanism to mitigate resentment, thereby improving human-AI cooperation.
-7.3. Future and Limitation
-This study offers valuable insights for AI agent designers and developers, emphasizing the importance of carefully considering target users, application contexts, and the purported characteristic of agents to meet practical requirements.
-The domain of AI-human collaborative interaction for task completion holds significant potential.
-Our research shifts the focus to competitive cooperation, addressing the growing prevalence of AI technologies capable of executing basic foundational tasks.
-While this represents a promising development, it also underscores the necessity of integrating more human-centered considerations into AI design.
-In future CSCW (Computer-Supported Cooperative Work) research, striking a balance between optimizing proclaimed characteristics and achieving cooperation efficiency remains an unresolved challenge.
-Although presenting an agent as a human can improve cooperative performance, it involves an element of deception, potentially leading to unforeseen risks if the agent’s true nature is disclosed.
-Additionally, designers may tailor an agent’s settings based on the gender of the human collaborator to enhance cooperation.
-However, it is crucial to ensure that such adjustments adhere to principles of transparency, fairness, and ethical responsibility.
-However, our study is limited by the controlled experimental settings, specifically the Prisoners’ Dilemma game, and the particular LLM agents employed.
-These controlled conditions may not fully capture the complexity and variability of real-world cooperative interactions.
-Additionally, a learning effect is inevitable when participants play the game consecutively, potentially altering their original behavioral strategies.
-Another imperfect aspect of our experiment is the lack of effective approaches to optimize the design and measurement methods for better comparing human-human and human-AI performance in competitive cooperation scenarios, which presents a challenge.
-To address these limitations, future studies should incorporate a broader range of LLM agents and more complex, realistic interaction scenarios.
-Based on this work, future research may explore the long-term effects of human interactions with LLM agents and examine the impact of continuous learning and adaptation in these agents.
-Given the potential widespread deployment of AI agents, policymakers should establish guidelines and ethical standards for their development and use.
-Ensuring that LLM agents’ behaviors and decision-making processes are reasonable is crucial for building public trust and acceptance.
-8. Conclusions
-With the continued proliferation of LLMs, these models are powered by an increasing number of applications and agents, driving a growing demand for deeper exploration of human-AI interaction.
-The transformation of AI from an assistant to an independent decision-making agent has already demonstrated significant potential and advantages in various contexts.
-It is crucial to investigate how humans are influenced by such agents, particularly in the domain of CSCW, where the dynamics of human-AI competitive cooperation merit further exploration.
-This study successfully verified the three hypotheses proposed at the outset of our study.
-Moreover, our research has identified distinct behavioral patterns in human interactions with LLM agents in competitive cooperation scenarios.
-Specifically, we observed differences in participants’ cooperation rates, decision-making times, amicability, and acceptance of repair efforts after betrayal when interacting with different purported-characteristic agents.
-These findings contribute to advancing the understanding of human-AI interaction by illuminating the nuanced ways in which humans adapt their behaviors based on their perceptions of various agents.
-It also addresses the relative lack of research on human-AI cooperation from a competitive perspective and provides experimental evidence of the value and potential of this line of inquiry.
-By uncovering the subtleties of human behavior in competitive cooperation, developers can design AI agents that promote improved collaboration and trust.
-More importantly, examining human cooperation with independent decision-making AI agents ahead of time offers critical insights and foresight for shaping ethical frameworks to guide future AI development.
+To optimize collaboration, embedding motivational scaffolding into AI architectures could transform competition into productive synergy.
+5.3. LLM Agents as Social Mirrors: Triggering Comparative Dynamics in Human Participants
+Our results indicate that, as referenced in Section 4.3, participants were more likely to experience the self-enhancement effect when interacting with LLM agents as opposed to purported rule-based AI agents (Alicke and Sedikides, 2009).
+We hypothesize that this is due to participants’ stronger perception of LLM agents’ capabilities, leading them to view these agents as more autonomous decision-makers.
+This, in turn, induces a stronger social presence effect (Oh et al., 2018), wherein LLM agents prompt participants to invest more time and effort in reflection and decision-making, thereby enhancing the depth and caution of their thinking.
+This is clearly reflected in the analysis of decision latency.
+In educational contexts, the integration of LLM agents can enhance student engagement and education performance (An et al., 2024; Song et al., 2024).
+As students interact with these agents, they may perceive them as more capable and intelligent than traditional educational tools, leading to greater investment in their learning.
+This heightened engagement can improve students’ ability to critically evaluate information and engage more deeply with content, much like the increased decision-making effort observed in our study when participants interacted with the LLM agent.
+While the social presence effect can enhance engagement, there are potential risks of over-reliance on LLM agents in education (Hamdi, 2024; Rizvi, 2023).
+If students view the LLM as an infallible authority, it may hinder their ability to think independently and critically evaluate information.
+This dependency could reduce the development of essential problem-solving skills, as students may rely on the AI for answers instead of engaging in deeper cognitive processes.
+Additionally, the social presence effect might cause stress or anxiety if students feel pressured to match the LLM’s capabilities, potentially affecting their mental well-being and creating an unhealthy learning environment.
+The comparative dynamics of LLM agents may worsen disparities in learning.
+Students struggling with the cognitive demands may feel inadequate, while those excelling may become overly reliant on the agent, distancing themselves from genuine understanding.
+This could lead to unequal learning outcomes, undermining the potential benefits of the LLM.
+Therefore, educators must strike a balance between the benefits of increased engagement and the need to foster independent thinking, ensuring that students learn to collaborate effectively with AI without becoming overly dependent on it.
+The goal is to help students develop the skills to solve complex problems while maintaining their ability to think independently.
+5.4. Limitation and Future
+While this study offers insights into human behavior in competitive cooperation with LLM agents, several limitations should be acknowledged.
+First, the study was conducted in a controlled experimental environment using a repeated Prisoner’s Dilemma task.
+While this allowed for systematic observation of decision-making patterns, such settings may not fully reflect the richness and unpredictability of real-world human-AI collaboration.
+Additionally, repeated interactions introduced potential learning effects, which may have influenced participants’ behavioral strategies over time.
+Second, the study focused on a limited set of agent types, with pre-defined “purported characteristics” such as being rule-based, human-like, or LLM-based.
+While these categories helped isolate specific effects, they may not capture the full diversity of AI systems or their evolving capabilities.
+Future studies could benefit from including a broader range of agent behaviors, learning capacities, and dynamic adaptation strategies to better approximate real-world deployment scenarios.
+Another challenge lies in the current lack of robust comparative frameworks for assessing human-human and human-AI cooperation.
+The development of more refined experimental designs and metrics—such as trust calibration, adaptive strategy shifts, and transparency evaluation—remains an open direction for future research.
+Furthermore, while our findings show that manipulating agent characteristics (e.g., by anthropomorphizing them or tailoring design based on gender) can influence cooperation outcomes, these approaches raise important ethical concerns.
+For example, portraying an agent as human may improve collaboration but introduces an element of deception, which could undermine user trust if revealed.
+Similarly, agent customization based on user demographics must be carefully examined to ensure fairness, transparency, and accountability.
+One promising avenue emerging from this study is the analysis of reflection time as a behavioral metric.
+Our results suggest that reflection time is sensitive to both the perceived identity of the agent and the participant’s own characteristics (e.g., gender).
+Longer reflection times were observed when participants believed they were competing with a human agent, and male participants deliberated longer when interacting with same-gender opponents.
+These findings align with prior literature on human complacency in AI-assisted decision-making and suggest that reflection time may serve as a useful proxy for cognitive engagement and trust calibration in human-AI interaction.
+Semi-structured interviews further revealed that participants often relied on prior assumptions or stereotypes when interpreting an agent’s behavior, which in turn shaped their responses.
+These insights highlight the need for future work to explore the cognitive and affective mechanisms underpinning human-agent interaction, especially in mixed-motive or competitive scenarios.
+In summary, while this study provides a foundation for understanding behavioral adaptation in competitive Human-AI cooperation, future work should explore more ecologically valid settings, refine evaluative frameworks, and consider the broader ethical implications of agent design and deployment.
+6. Conclusion
+As large language models (LLMs) evolve beyond assistive roles to take on autonomous decision-making, new challenges emerge for human-AI interaction, especially in contexts involving both collaboration and competition.
+This study addressed three research questions concerning human behavior in competitive cooperation with LLM agents.
+Our findings revealed that participants’ cooperation rates, decision latency, unsolicited cooperation, and trust restoration varied depending on the perceived characteristics of the AI agents and the human’s gender.
+These differences suggest that users adapt their strategies based on how they interpret an agent’s intent and autonomy.
+By examining mixed-motive interactions, this study contributes to a growing body of work that moves beyond traditional assistive paradigms.
+It offers empirical insights into how humans adjust their behavior when facing agentic AI systems, highlighting the need for context-sensitive and ethically grounded design.
+These findings underscore the importance of anticipating user expectations and perceptions as AI systems increasingly participate in foundational tasks, offering a timely foundation for responsible AI development and deployment.